### v1

# Cartesian subgroups in graph products of groups

Abstract.
The kernel of the natural projection of a graph product of groups onto their direct product is called the Cartesian subgroup. This construction generalises commutator subgroups of right-angled Coxeter and Artin groups. Using theory of polyhedral products, we give a lower and an upper bound on the number of relations in presentations of Cartesian groups and on their deficiency. The bounds are related to the fundamental groups of full subcomplexes in the clique complex, and the lower bound coincide with the upper bound if these fundamental groups are free or free abelian.
Following Li Cai’s approach, we also describe an algorithm that computes “small” presentations of Cartesian subgroups.
Key words and phrases:
commutator subgroup, right-angled Coxeter group, graph product, polyhedral product
2020 Mathematics Subject Classification:
20F05, 20F55, 57M07; 20F36, 20F65, 57M05, 57S12
1. Introduction
Let $\Gamma$ be a simple graph on the vertex set $[m]=\{1,\dots,m\}$ and $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of discrete groups. The corresponding graph product [Gre90]
$\underline{G}^{\Gamma}:=(G_{1}\ast\dots\ast G_{m})/(g_{i}g_{j}=g_{j}g_{i},~{}\forall\{i,j\}\in\Gamma,~{}\forall g_{i}\in G_{i},~{}\forall g_{j}\in G_{j})$
interpolates between the free product and the direct product of $G_{1},\dots,G_{m}$ as $\Gamma$ varies between the edgeless graph and the complete graph. The Cartesian subgroup
$\operatorname{Cart}(\underline{G},\Gamma):=\operatorname{Ker}(\underline{G}^{\Gamma}\to G_{i}\times\dots\times G_{m})$
[…]
The right-angled Coxeter groups
$\mathrm{RC}_{\Gamma}:=\langle g_{1},\dots,g_{m}\mid g_{i}^{2}=1,~{}i=1,\dots,m;~{}g_{i}g_{j}=g_{j}g_{i},~{}\{i,j\}\in\Gamma\rangle=(\underline{\mathbb{Z}_{2}})^{\Gamma}$
and right-angled Artin groups $\mathrm{RA}_{\Gamma}:=\langle g_{1},\dots,g_{m}\mid g_{i}g_{j}=g_{j}g_{i},~{}\{i,j\}\in\Gamma\rangle=\underline{\mathbb{Z}}^{\Gamma}$ are special cases of the graph product construction. Their Cartesian subgroups are the commutator subgroups $\mathrm{RC}_{\Gamma}^{\prime}=\operatorname{Cart}(\underline{\mathbb{Z}_{2}},\Gamma)$ and $\mathrm{RA}_{\Gamma}^{\prime}=\operatorname{Cart}(\underline{\mathbb{Z}},\Gamma).$ The classifying spaces of graph products and their Cartesian subgroups can be described as certain polyhedral products of topological spaces [PV16, BBC19] (see Proposition 2.5); conversely, the fundamental groups of some polyhedral products can be identified with $\underline{G}^{\Gamma}$ or $\operatorname{Cart}(\underline{G},\Gamma)$. For example, each simplicial complex $\mathcal{K}$ corresponds to the real moment-angle complex $\mathcal{R}_{\mathcal{K}}$, a polyhedral product important in toric topology [BP15]. It is known that $\pi_{1}(\mathcal{R}_{\mathcal{K}})\cong\mathrm{RC}^{\prime}_{\Gamma}=\operatorname{Cart}(\underline{\mathbb{Z}_{2}},\Gamma)$ where $\Gamma=\operatorname{sk}_{1}\mathcal{K}$ is the $1$-skeleton of $\mathcal{K}$. (In general, fundamental groups of polyhedral products are relative graph products, see [Dav12, Theorem 2.18].)
We study presentations of Cartesian subgroups and related numerical invariants: their rank and deficiency
$\operatorname{rank}G:=\inf_{G=\langle X\mid R\rangle}|X|,\quad\operatorname{def}G:=\inf_{G=\langle X\mid R\rangle}(|R|-|X|).$
It is convenient to consider the clique complex of the graph $\Gamma$
$\mathcal{K}(\Gamma):=\{I\subset[m]:~{}\{i,j\}\in\Gamma,~{}\forall i,j\in I\},$
the unique flag simplicial complex $\mathcal{K}$ such that $\Gamma=\operatorname{sk}_{1}\mathcal{K}.$ From now on we assume that $\mathcal{K}=\mathcal{K}(\Gamma)$ and write $\underline{G}^{\mathcal{K}}:=\underline{G}^{\Gamma},$ etc, by abuse of notation.
The ranks of Cartesian subgroups were computed by Panov and Veryovkin [PV19, Theorem 5.2(b)]: $\operatorname{rank}\operatorname{Cart}(\underline{G},\mathcal{K})=N(\underline{G},\mathcal{K}),$ where
$N(\underline{G},\mathcal{K}):=\sum_{J\subset[m]}n_{J}\cdot\widetilde{b}_{0}(\mathcal{K}_{J}),$
$n_{J}:=\prod_{j\in J}|G_{j}\setminus\{1_{j}\}|$ for $J\subset[m]$ and $\widetilde{b}_{i}(X):=\dim_{\mathbb{Q}}{\widetilde{H}}_{i}(X;\mathbb{Q}).$ (In particular, the induced subgraph $\Gamma_{J}$ of $\Gamma$ has $\widetilde{b}_{0}(\mathcal{K}_{J})+1$ path components.) Moreover, Panov and Veryovkin described an explicit minimal generating set for $\operatorname{Cart}(\underline{G},\mathcal{K})$ that consists of nested iterated commutators [PV19, Theorem 5.2(a)]. A special case of this result, the formula $\operatorname{rank}\mathrm{RC}_{\mathcal{K}}^{\prime}=\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})$ and a set of $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})$ generators for the group $\mathrm{RC}_{\mathcal{K}}^{\prime}$, was obtained earlier by the same authors [PV16, Theorem 4.5].
[…]
Theorem 1.1.
Let $\mathcal{K}$ be a flag simplicial complex on the vertex set $[m]$ and $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of discrete groups. Then the group $\operatorname{Cart}(\underline{G},\mathcal{K})=\operatorname{Ker}(\underline{G}^{\mathcal{K}}\to\prod_{i=1}^{m}G_{i})$ admits a presentation by $N(\underline{G},\mathcal{K})=\sum_{J\subset[m]}n_{J}\cdot\widetilde{b}_{0}(\mathcal{K}_{J})$ generators modulo
$M^{+}(\underline{G},\mathcal{K}):=\sum_{J\subset[m]}n_{J}\cdot\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$
relations, where $n_{J}:=\prod_{j\in J}|G_{j}\setminus\{1_{j}\}|$. In particular,
$\operatorname{def}\operatorname{Cart}(\underline{G},\mathcal{K})\leq M^{+}(\underline{G},\mathcal{K})-N(\underline{G},\mathcal{K}).$
The generators are explicit (each generator is of the form $L_{\underline{g}}(i,J):=\prod_{j\in J}g_{j}\cdot g_{i}^{-1}\cdot(\prod_{j\in J\setminus i}g_{j})^{-1}$), while the relations can be computed by a recursive algorithm (see Theorem 4.11). Moreover, the generators and relations have a geometric meaning in terms of paths in certain polyhedral products, and this result can be extended to the case of infinite graph products (see Subsection 6.1). We also give a homological lower bound on the size of a presentation:
Theorem 1.2.
Any presentation of the group $\operatorname{Cart}(\underline{G},\mathcal{K})$ contains at least
$M^{-}(\underline{G},\mathcal{K}):=\operatorname{rank}\bigoplus_{J\subset[m]}{\widetilde{H}}_{1}(\mathcal{K}_{J};\mathbb{Z})^{\oplus n_{J}}$
relations. Moreover, $\operatorname{def}\operatorname{Cart}(\underline{G},\mathcal{K})\geq M^{-}(\underline{G},\mathcal{K})-N(\underline{G},\mathcal{K}).$
Since $(\Pi_{1}(X))_{\operatorname{ab}}={\widetilde{H}}_{1}(X;\mathbb{Z})$, the numbers $M^{-}(\underline{G},\mathcal{K})$ and $M^{+}(\underline{G},\mathcal{K})$ coincide if all full subcomplexes of $\mathcal{K}$ have free or free abelian fundamental groups. This is the case when $m$ is small enough (see Remark 6.1) or if $\mathcal{K}$ is a part of a surface triangulation (see Proposition 6.3). In these cases, we know both the rank and deficiency of Cartesian subgroups.
As a special case we obtain a small presentation of the commutator subgroup $\mathrm{RC}_{\mathcal{K}}^{\prime}$ in any right-angled Coxeter group, and a lower bound on the number of relations in any its presentation:
Corollary 1.3.
Let $\mathcal{K}$ be a flag simplicial complex on the vertex set $[m]$. Then
  1. (1)
    The commutator subgroup $\mathrm{RC}_{\mathcal{K}}^{\prime}$ of the corresponding right-angled Coxeter group $\mathrm{RC}_{\mathcal{K}}$ admits a presentation by $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})$ generators modulo $\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$ relations.
  2. (2)
    Any presentation of the group $\mathrm{RC}_{\mathcal{K}}^{\prime}$ contains at least
    $\operatorname{rank}\bigoplus_{J\subset[m]}{\widetilde{H}}_{1}(\mathcal{K}_{J};\mathbb{Z})$ relations.
  3. (3)
    The deficiency $\operatorname{def}\mathrm{RC}_{\mathcal{K}}^{\prime}$ satisfies the inequalities
    $\operatorname{rank}\bigoplus_{J\subset[m]}{\widetilde{H}}_{1}(\mathcal{K}_{J};\mathbb{Z})\leq\operatorname{def}\mathrm{RC}_{\mathcal{K}}^{\prime}+\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\leq\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J}).\qed$
Two earlier results can be deduced from the theorems above:
  • Characterisation of simplicial complexes $\mathcal{K}$ such that $\operatorname{Cart}(\underline{G},\mathcal{K})$ is a free group [HR12, Theorem 3.2] (independently obtained in [PV16, Theorem 4.3]);
  • Characterisation of simplicial complexes $\mathcal{K}$ such that $\mathrm{RC}_{\mathcal{K}}^{\prime}$ is a one-relator group [GIPS22, Theorem 3.2(a)].
Structure of paper
In Section 2 we discuss necessary definitions and known results about simplicial complexes, graph products of groups and polyhedral products of topological spaces. In Section 3 we prove the lower bound (Theorem 1.2). In Section 4 we describe an explicit set of generators for $\operatorname{Cart}(\underline{G},\mathcal{K})$ and an algorithm that computes a presentation on these generators (see Theorem 4.11 which implies Theorem 1.1). In Section 5 we prove Theorem 4.11, applying a version of the van Kampen theorem to a certain explicit classifying space of $\operatorname{Cart}(\underline{G},\mathcal{K})$. Finally, in section 6 we discuss possible generalisations and discuss the similarities between results of this paper and known results on graph products of algebras.
Acknowledgements and funding
This work was carried out within the project “Mirror Laboratories” of HSE University, Russian Federation, and is supported by the Russian Science Foundation under grant no. 24-71-00059, http://rscf.ru/en/project/24-71-00059/.
The author woud like to thank Taras E. Panov for suggesting the problem and valuable advice, and Li Cai, Anton A. Klyachko, Temurbek Rahmatullaev and Andrei Yu. Vesnin for helpful discussions.
2. Preliminaries
2.1. Simplicial complexes
A simplicial complex $\mathcal{K}$ on the vertex set $V$ is a non-empty collection of finite subsets $I\subset V,$ called the faces, that is closed under inclusion (i.e. if $J\subset I\in\mathcal{K}$, then $J\in\mathcal{K}$). Usually $V=[m]:=\{1,\dots,m\}.$ If $i\in V$ and $\{i\}\notin\mathcal{K}$, then $i$ is called a ghost vertex. We consider only complexes without ghost vertices, so we assume that $\{i\}\in\mathcal{K}$ for every $i\in V.$
Given $J\subset V,$ the simplicial complex $\mathcal{K}_{J}:=\{I\in\mathcal{K}:~{}I\subset J\}$ on vertex set $J$ is called a full subcomplex of $\mathcal{K}.$
A missing face of $\mathcal{K}$ is a set $J\subset V$ such that $J\notin\mathcal{K}$, but every proper subset of $J$ is a face of $\mathcal{K}.$ A complex $\mathcal{K}$ is flag if all of its missing faces consist of two vertices; equivalently, a flag complex is the clique complex of its 1-skeleton ($I\in\mathcal{K}$ if and only if $\{i,j\}\in\operatorname{sk}_{1}\mathcal{K}$ for all $i,j\in I$). Every full subcomplex of a flag complex is flag.
Every face $I\in\mathcal{K}$ corresponds to a geometric face $|I|=\mathrm{conv}\{e_{i}:~{}i\in I\}$ of the standard geometric simplex $L^{m-1}=\mathrm{conv}\{e_{i}:i\in[m]\}\subset\mathbb{R}^{m}.$ The union of these geometric faces is called the geometric realization $|\mathcal{K}|$ of a complex $\mathcal{K}.$
2.2. Polyhedral products
Let $(\underline{X},\underline{A})=\{(X_{i},A_{i})\}_{i=1}^{m}$ be a sequence of pairs of topological spaces and $\mathcal{K}$ be a simplicial complex on $[m].$ The corresponding polyhedral product is the following subspace of $X_{1}\times\dots\times X_{m}$:
$(\underline{X},\underline{A})^{\mathcal{K}}:=\bigcup_{I\in\mathcal{K}}\Big{(}\prod_{i\in I}X_{i}\times\prod_{i\in[m]\setminus I}A_{i}\Big{)}.$
It is clear that $(\underline{X},\underline{A})^{\mathcal{K}}\simeq(\underline{Y},\underline{B})^{\mathcal{K}}$ if $(X_{i},A_{i})\simeq(Y_{i},B_{i})$ for all $i=1,\dots,m$ (the symbol $\simeq$ denotes homotopy equivalence). We write $(X,A)^{\mathcal{K}}:=(\underline{X},\underline{A})^{\mathcal{K}}$ if $X_{1}=\dots=X_{m}=X,$ $A_{1}=\dots=A_{m}=A.$ For a sequence of pointed spaces $(X_{1},\dots,X_{m})$ denote also $\underline{X}^{\mathcal{K}}:=(\underline{X},\underline{{\mathrm{pt}}})^{\mathcal{K}}.$
The real moment-angle complex $\mathcal{R}_{\mathcal{K}}:=(D^{1},S^{0})^{\mathcal{K}}$ is a special case of this construction. Homology groups of real moment-angle complexes are well known.
Proposition 2.1 (see [BP15, Theorem 4.5.8]).
$H_{i}(\mathcal{R}_{\mathcal{K}})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J}).$
A far-reaching generalization of this result is the following stable homotopy decomposition for polyhedral products of the form $(\operatorname{cone}\underline{A},\underline{A})^{\mathcal{K}}:$
Theorem 2.2 ([BBCG10, Theorem 2.21]).
Let $(\underline{X},\underline{A})=\{(X_{i},A_{i})\}_{i=1}^{m}$ be a sequence of CW pairs such that all $X_{i}$ are contractible. Then there is a homotopy equivalence
$\Sigma(\underline{X},\underline{A})^{\mathcal{K}}\simeq\Sigma^{2}\bigvee_{J\subset[m]}|\mathcal{K}_{J}|\wedge\underline{A}^{\wedge J},$
where $\underline{A}^{\wedge J}:=\bigwedge_{j\in J}A_{j}$ is the smash product.∎
Corollary 2.3.
Under the conditions of Theorem 2.2, we have
${\widetilde{H}}_{i}((\underline{X},\underline{A})^{\mathcal{K}};\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(|\mathcal{K}_{J}|\wedge\underline{A}^{\wedge J};\mathbb{Z}).\qed$
2.3. Classifying spaces of graph products and Cartesian subgroups
For a topological group $G$, there is a principal $G$-fibration $EG\to BG$ with $EG$ contractible, which is unique up to a weak homotopy equivalence. The base space $BG$ is called the classifying space of $G.$ For discrete groups the classifying space $BG=K(G,1)$ is aspherical and $EG$ is its universal cover. In the next proposition we denote $E\underline{G}:=(EG_{1},\dots,EG_{m})$ and $B\underline{G}:=(BG_{1},\dots,BG_{m}).$
Proposition 2.4 (see e.g. [PV16, Proposition 3.1]).
Let $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of topological groups and $\mathcal{K}$ be a simplicial complex on $[m]$. Then there is a canonical homotopy fibration $(E\underline{G},\underline{G})^{\mathcal{K}}\to(B\underline{G})^{\mathcal{K}}\to\prod_{i=1}^{m}BG_{i}.\qed$
Combining this fibration with the results of Panov, Ray and Vogt [PRV04, Proposition 5.1] provides an important connection between polyhedral products of spaces and graph products of discrete groups:
Proposition 2.5 ([PV16, Theorem 3.2]).
Let $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of discrete groups and $\mathcal{K}$ be a simplicial complex on $[m]$. Denote $\Gamma=\operatorname{sk}_{1}\mathcal{K}$. Then
  1. (1)
    $\pi_{1}((B\underline{G})^{\mathcal{K}})\cong\underline{G}^{\Gamma}$ and $\pi_{1}((E\underline{G},\underline{G})^{\mathcal{K}})\cong\operatorname{Cart}(\underline{G},\Gamma)$;
  2. (2)
    $\pi_{k}((B\underline{G})^{\mathcal{K}})\cong\pi_{k}((E\underline{G},\underline{G})^{\mathcal{K}}),~{}k\geq 2;$
  3. (3)
    $(E\underline{G},\underline{G})^{\mathcal{K}}$ and $(B\underline{G})^{\mathcal{K}}$ are aspherical if and only if $\mathcal{K}$ is a flag complex.
In particular, for flag complexes $\mathcal{K}$ we have
$B(\underline{G}^{\mathcal{K}})=(B\underline{G})^{\mathcal{K}},\quad B(\operatorname{Cart}(\underline{G},\mathcal{K}))=(E\underline{G},\underline{G})^{\mathcal{K}}.\qed$
Remark 2.6.
Davis described the fundamental groups $\pi_{1}((\underline{X},\underline{A})^{\mathcal{K}})$ of general polyhedral products algebraically as relative graph products [Dav12, Theorem 2.18].
Asphericity of $(E\underline{G},\underline{G})^{\mathcal{K}}$ in the flag case can also be proved by applying the theory of CAT(0)-spaces to the universal covering over the cubical complex $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, see [Dav08, Chapter 12] for $G_{i}=\mathbb{Z}_{2}$ case and [Dav12, Lemma 2.11] in the general case. This approach can be traced back to Meier [Mei96, Section 4].
An almost complete answer to the question “Which polyhedral products are aspherical?” was obtained by Davis and Kropholler [DK15, Theorem 1].
Recall that the right-angled Coxeter group
$\mathrm{RC}_{\mathcal{K}}:=\langle g_{1},\dots,g_{m}\mid g_{i}^{2}=1,~{}i=1,\dots,m;~{}g_{i}g_{j}=g_{j}g_{i},~{}\{i,j\}\in\mathcal{K}\rangle$
is the graph product of groups $G_{1}=\dots=G_{m}=\mathbb{Z}_{2}.$
Corollary 2.7 ([PV16, Corollary 3.4]).
Let $\mathcal{K}$ be a flag simplicial complex. Then
$B(\mathrm{RC}_{\mathcal{K}})=(\mathbb{R}P^{\infty})^{\mathcal{K}},\quad B(\mathrm{RC}_{\mathcal{K}}^{\prime})=\mathcal{R}_{\mathcal{K}}.$
Proof.
Indeed, by Proposition 2.5 we have $B(\mathrm{RC}_{\mathcal{K}})=(B\mathbb{Z}_{2})^{\mathcal{K}}=(\mathbb{R}P^{\infty})^{\mathcal{K}}$ and $B(\mathrm{RC}^{\prime}_{\mathcal{K}})=(E\mathbb{Z}_{2},\mathbb{Z}_{2})^{\mathcal{K}}=(S^{\infty},S^{0})^{\mathcal{K}}\simeq(D^{1},S^{0})^{\mathcal{K}}=\mathcal{R}_{\mathcal{K}}$. ∎
The isomorphism $\pi_{1}(\mathcal{R}_{\mathcal{K}})\cong\mathrm{RC}_{\mathcal{K}}^{\prime}$ has a geometric interpretation: each generator $g_{i}\in\mathrm{RC}_{\mathcal{K}}$ corresponds to the path along the $i$th coordinate edge of the cube $(D^{1})^{m}\supset\mathcal{R}_{\mathcal{K}}$. Group multiplication corresponds to concatenation of paths. Then the word $w\in\mathrm{RC}_{\mathcal{K}}$ belongs to the commutator subgroup if and only if the corresponding path is a loop. Every edge $\{i,j\}\in\mathcal{K}$ corresponds, from the algebraic point of view, to a relation $g_{i}g_{j}=g_{j}g_{i}$ in $\mathrm{RC}_{\mathcal{K}},$ and, from the topological point of view, to $2^{m-2}$ squares in $\mathcal{R}_{\mathcal{K}}.$ Due to the presence of these squares, the paths $\dots g_{i}g_{j}\dots$ and $\dots g_{j}g_{i}\dots$ are homotopic in $\mathcal{R}_{\mathcal{K}}.$ Thus a homomorphism $\mathrm{RC}_{\mathcal{K}}^{\prime}\to\pi_{1}(\mathcal{R}_{\mathcal{K}})$ is well defined; in fact, it is an isomorophism. Its generalisation $\operatorname{Cart}(\underline{G},\mathcal{K})\overset{\cong}{\longrightarrow}\pi_{1}((\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}})$ is discussed in Subsection 5.5 below.
2.4. Ranks of groups and free products
The rank of a group $G$ is the smallest number $N$ such that $G$ can be generated by $N$ elements. (If $G$ is not finitely generated, $N$ is considered as a cardinal number.)
Proposition 2.8.
The rank has the following properties:
  1. (1)
[…]
  2. (2)
    If $f:G\to H$ is a surjective homomorphism, then $\operatorname{rank}G\geq\operatorname{rank}H$;
  3. (3)
    $\operatorname{rank}\mathbb{Z}^{m}=m.$
Proof.
  1. (1)
    This is [LS77, Theorem 1.8].
  2. (2)
    If $G$ is generated by $g_{1},\dots,g_{N},$ then $H$ is generated by $f(g_{1}),\dots,f(g_{N}).$
  3. (3)
    If $\mathbb{Z}^{m}$ could be generated by $n<m$ elements, then the vector space $\mathbb{Q}^{m}$ would be a linear span of $n<m$ elements. ∎
Definition 2.9.
[…]
Proposition 2.10.
Let $X$ be a topological space and $\{X_{\alpha}\}$ be the set of its path components. Then
  1. (1)
    $\operatorname{rank}\Pi_{1}(X)=\sum_{\alpha}\operatorname{rank}\pi_{1}(X_{\alpha})$;
  2. (2)
    $\Pi_{1}(X)_{\operatorname{ab}}\cong H_{1}(X;\mathbb{Z})$;
  3. (3)
    $\Pi_{1}(X)\cong\pi_{1}(\bigvee_{\alpha}X_{\alpha})$.
Proof.
Statement (1) follows from the Grushko’s theorem, (2) from the Poincaré–Hurewicz theorem [Hat02, Theorem 2A.1] and (3) from the van Kampen theorem [Hat02, Theorem 1.20]. ∎
3. Proof of Theorem 1.2
3.1. A general lower bound on the number of generators and relations
Recall that the integer homology of a discrete group is isomorphic to the integer homology of its classifying space [Bro82, Proposition II.4.1]:
$\mathrm{H}_{i}(G;\mathbb{Z}):=\operatorname{Tor}^{\mathbb{Z}[G]}_{i}(\mathbb{Z},\mathbb{Z})\cong H_{i}(BG;\mathbb{Z}).$
The following lemma is essentially [Bro82, §II.5, Exercise 5a]. Notably, it has a purely algebraic proof [Eps61, Lemma 1.2].
Lemma 3.1.
Let $G=\langle x_{1},\dots,x_{N}\mid r_{1},\dots,r_{M}\rangle$ be a finite group presentation. Then
$N\geq\operatorname{rank}\mathrm{H}_{1}(G;\mathbb{Z}),\quad M-N\geq\operatorname{rank}\mathrm{H}_{2}(G;\mathbb{Z})-\dim_{\mathbb{Q}}\mathrm{H}_{1}(G;\mathbb{Q}).$
In particular, $M\geq\operatorname{rank}\mathrm{H}_{2}(G;\mathbb{Z}).$
Proof.
We construct a CW classifying space for the group $G$ by attaching cells to its presentation complex. Then $BG=\bigvee_{i=1}^{N}S_{i}^{1}\cup\bigcup_{j=1}^{M}e_{j}^{2}\cup\bigcup_{\beta}e_{\beta}^{>2},$ where the 2-cells are attached by the maps corresponding to the words $r_{1},\dots,r_{M}\in F(x_{1},\dots,x_{N})\cong\pi_{1}(\bigvee_{i=1}^{N}S_{i}^{1}).$ The cellular chain complex of $BG$ is of the form
$0\leftarrow\mathbb{Z}\overset{~{}0}{\longleftarrow}\mathbb{Z}^{N}\overset{~{}\partial_{2}}{\longleftarrow}\mathbb{Z}^{M}\overset{~{}\partial_{3}}{\longleftarrow}\dots$
and has $H_{*}(BG;\mathbb{Z})\cong\mathrm{H}_{*}(G;\mathbb{Z})$ as its homology.
Denote $k=\operatorname{rank}\partial_{2}$. Then $\mathrm{H}_{1}(G;\mathbb{Z})$ is a quotient of $\mathbb{Z}^{N}$, $\mathrm{H}_{2}(G;\mathbb{Z})$ is a quotient of $\operatorname{Ker}\partial_{2}\simeq\mathbb{Z}^{M-k}$, and $\mathrm{H}_{1}(G;\mathbb{Q})\simeq\mathbb{Q}^{N-k}$. Hence $\operatorname{rank}\mathrm{H}_{1}(G;\mathbb{Z})\leq N$ and $\operatorname{rank}\mathrm{H}_{2}(G;\mathbb{Z})-\dim_{\mathbb{Q}}\mathrm{H}_{1}(G;\mathbb{Q})\leq(M-k)-(N-k)=M-N$. ∎
Proposition 3.2.
Let $G=\langle x_{1},\dots,x_{N}\mid r_{1},\dots,r_{M}\rangle$ be a finitely presented group such that $\mathrm{H}_{1}(G;\mathbb{Z})\simeq\mathbb{Z}^{N}.$ Then $r_{1},\dots,r_{M}$ belong to the commutator subgroup of $F(x_{1},\dots,x_{N}).$
Proof.
In the notation of the previous lemma, we have $\mathbb{Z}^{N}/\operatorname{Im}\partial_{2}\simeq\mathbb{Z}^{N},$ hence $\partial_{2}=0.$ On the other hand, $\partial_{2}:\mathbb{Z}^{M}\to\mathbb{Z}^{N}$ is the abelianization of the map $F(r_{1},\dots,r_{M})\to F(x_{1},\dots,x_{N}).$ Therefore $r_{j}\in\operatorname{Ker}(F(x_{1},\dots,x_{N})\overset{\operatorname{ab}}{\longrightarrow}\mathbb{Z}^{N})=F(x_{1},\dots,x_{N})^{\prime}.$
Remark 3.3.
For a finitely generated abelian group $A,$ let $\mathrm{gen}\,A$ and $\mathrm{rel}\,A$ be the smallest numbers such that $A\simeq\mathbb{Z}^{\mathrm{gen}\,A}/\mathbb{Z}^{\mathrm{rel}\,A}.$ Clearly, $\mathrm{gen}\,A=\operatorname{rank}A$ and $\mathrm{rel}\,A=\operatorname{rank}A-\dim_{\mathbb{Q}}(A\otimes_{\mathbb{Z}}\mathbb{Q}).$ Hence Lemma 3.1 can be stated as
$N\geq\mathrm{gen}\,\mathrm{H}_{1}(G;\mathbb{Z}),\quad M-N\geq\mathrm{gen}\,\mathrm{H}_{2}(G;\mathbb{Z})-\mathrm{gen}\,\mathrm{H}_{1}(G;\mathbb{Z})+\mathrm{rel}\,\mathrm{H}_{1}(G;\mathbb{Z}).$
Since $\mathrm{H}_{i}(G;\mathbb{Z})=\operatorname{Tor}^{\mathbb{Z}[G]}_{i}(\mathbb{Z},\mathbb{Z}),$ we deduce: if $G$ has a presentation by $N$ generators and $M$ relations, then
$N\geq\mathrm{gen}\,\operatorname{Tor}^{\mathbb{Z}[G]}_{1}(\mathbb{Z},\mathbb{Z}),~{}M\geq\mathrm{gen}\,\operatorname{Tor}^{\mathbb{Z}[G]}_{2}(\mathbb{Z},\mathbb{Z})+\mathrm{rel}\,\operatorname{Tor}^{\mathbb{Z}[G]}_{1}(\mathbb{Z},\mathbb{Z}).$
There is a similar result on presentations of connected graded associative algebras with unit [Vyl24, Theorem A.10]: if $\mathbf{k}$ is a principal ideal domain and $R=\bigoplus_{n\geq 0}R_{n}$ is a connected graded $\mathbf{k}$-algebra presented by $N$ generators modulo $M$ relations, then
$N\geq\mathrm{gen}\,\operatorname{Tor}^{R}_{1}(\mathbf{k},\mathbf{k}),~{}M\geq\mathrm{gen}\,\operatorname{Tor}^{R}_{2}(\mathbf{k},\mathbf{k})+\mathrm{rel}\,\operatorname{Tor}^{R}_{1}(\mathbf{k},\mathbf{k}).$
Moreover, there is a “minimal” presentation of $R$ such that both lower bounds are achieved. This is not true for groups: if $G$ is a knot group, then $\mathrm{H}_{1}(G;\mathbb{Z})=\mathbb{Z},$ but usually $\operatorname{rank}G>1$.
3.2. Homology of Cartesian subgroups
Proposition 3.4.
Let $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of discrete groups and $\mathcal{K}$ be a flag simplicial complex on vertex set $[m]$. Then
$\mathrm{H}_{i}(\operatorname{Cart}(\underline{G},\mathcal{K});\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J};\mathbb{Z})^{\oplus n_{J}},$
where $n_{J}:=\prod_{j\in J}|G_{j}\setminus\{1_{j}\}|.$ In particular, $\mathrm{H}_{i}(\mathrm{RC}_{\mathcal{K}}^{\prime};\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J};\mathbb{Z}).$
Proof.
We have $\mathrm{H}_{*}(\operatorname{Cart}(\underline{G},\mathcal{K});\mathbb{Z})=H_{*}((E\underline{G},\underline{G})^{\mathcal{K}};\mathbb{Z})$ by Proposition 2.5, and
$H_{i}((E\underline{G},\underline{G})^{\mathcal{K}};\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(|\mathcal{K}_{J}|\wedge\underline{G}^{\wedge J};\mathbb{Z})$
by Corollary 2.3. Note that $G_{j}$ are discrete spaces. If $A$ and $B$ are discrete and $|A|=n+1,$ $|B|=m+1$ then $|A\wedge B|$ is discrete and has $mn+1$ points; hence $\underline{G}^{\wedge J}$ is discrete and has $n_{J}+1$ points. Finally, if $A$ is discrete and $|A|=n+1,$ then $X\wedge A\cong X^{\vee n},$ hence ${\widetilde{H}}_{*}(X\wedge A;\mathbb{Z})\cong{\widetilde{H}}_{*}(X;\mathbb{Z})^{\oplus n}.$
Proof of Theorem 1.2.
We have
$\operatorname{rank}\mathrm{H}_{1}(\operatorname{Cart}(\underline{G},\mathcal{K});\mathbb{Z})=N(\underline{G},\mathcal{K}),\quad\operatorname{rank}\mathrm{H}_{2}(\operatorname{Cart}(\underline{G},\mathcal{K});\mathbb{Z})=M^{-}(\underline{G},\mathcal{K})$
by Proposition 3.4. Then, by Lemma 3.1, the deficiency of $\operatorname{Cart}(\underline{G},\mathcal{K})$ is not less than $M^{-}(\underline{G},\mathcal{K})-N(\underline{G},\mathcal{K})$, and any presentation of this group has at least $M^{-}(\underline{G},\mathcal{K})$ relations. ∎
We also obtain a curious property of “minimal presentations” of $\operatorname{Cart}(\underline{G},\mathcal{K})$, which is not obvious for the presentation that will be obtained in Theorem 4.11.
Proposition 3.5.
Suppose that the number $N(\underline{G},\mathcal{K})$ is finite, and let $\operatorname{Cart}(\underline{G},\mathcal{K})=\langle x_{1},\dots,x_{N(\underline{G},\mathcal{K})}\mid R_{1},\dots,R_{M}\rangle$ be a group presentation. Then the relations $R_{1},\dots,R_{M}$ belong to the commutator subgroup of $F(x_{1},\dots,x_{N(\underline{G},\mathcal{K})}).$
Proof.
This follows from Proposition 3.2, since $\mathrm{H}_{1}(\operatorname{Cart}(\underline{G},\mathcal{K});\mathbb{Z})\simeq\mathbb{Z}^{N(\underline{G},\mathcal{K})}.$
4. Explicit presentations of Cartesian subgroups in graph products
In Theorem 4.11 we will describe a small presentation of the group $\operatorname{Cart}(\underline{G},\mathcal{K})$, clarifying Theorem 1.1. The proof will be given in Section 5.
Denote by $F(X)$ the free group generated by a set $X$. We first describe a set of elements $D=\{L_{\underline{g}}(i,J):\dots\}\subset\operatorname{Cart}(\underline{G},\mathcal{K})$ and its subset $\widehat{D}\subset D$ of distinguished elements. Then we show that each element of $D$ is equal in $\operatorname{Cart}(\underline{G},\mathcal{K})$ to a word on the distinguished elements, and provide an algorithm that computes such words. Hence each word $r\in F(D)$ is equal in $\operatorname{Cart}(\underline{G},\mathcal{K})$ to a word $\operatorname{Red}(r)\in F(\widehat{D})$ which can be computed inductively. Our presentation is of the form
$\operatorname{Cart}(\underline{G},\mathcal{K})=\langle\widehat{D}\mid\operatorname{Red}(R_{\underline{g}}(\lambda,J)):\dots\rangle,$
where the words $R_{\underline{g}}(\lambda,J)\in F(\widehat{D})$ are known explicitly.
4.1. Generators
Fix a sequence $\underline{G}=(G_{1},\dots,G_{m})$ of discrete groups and a flag simplicial complex $\mathcal{K}$ on $[m]$. Denote $G_{j}^{*}:=G_{j}\setminus\{1_{j}\}$.
Definition 4.1.
For $i\in J\subset[m]$ and ${\underline{g}}=(g_{j}:j\in J)\in\prod_{j\in J}G_{j}^{*},$ define
$L_{\underline{g}}(i,J):=\prod_{j\in J}g_{j}\cdot g_{i}^{-1}\cdot\Big{(}\prod_{j\in J\setminus i}g_{j}\Big{)}^{-1}\in\operatorname{Cart}(\underline{G},\mathcal{K}),$
where the products are in the ascending order. Denote
$D:=\Big{\{}L_{\underline{g}}(i,J):J\subset[m],~{}i\in J,~{}{\underline{g}}\in\prod_{j\in J}G_{j}^{*}\Big{\}}\subset\operatorname{Cart}(\underline{G},\mathcal{K}).$
For example, $L_{\underline{g}}(3,\{1,2,3,8\})=g_{1}g_{2}g_{3}g_{8}\cdot g_{3}^{-1}\cdot g_{8}^{-1}g_{2}^{-1}g_{1}^{-1}.$
We use the following notation from [Vyl24, Definition 5.2] inspired by [GPTW16].
Definition 4.2.
For each $J\subset[m]$, choose a subset $\Theta(J)\subset J\setminus\{\max(J)\}$ such that $\Theta(J)\sqcup\{\max(J)\}$ contains exactly one vertex from each path component of $\mathcal{K}_{J}$.
We define the set of distinguished elements
$\widehat{D}:=\Big{\{}L_{\underline{g}}(i,J):J\subset[m],~{}i\in\Theta(J),~{}{\underline{g}}\in\prod_{j\in J}G_{j}^{*}\Big{\}}\subset D.$
Since $|\Theta(J)|=\widetilde{b}_{0}(\mathcal{K}_{J})=b_{0}(\mathcal{K}_{J})-1$, we have $|\widehat{D}|=\sum_{J\subset[m]}n_{J}\widetilde{b}_{0}(\mathcal{K}_{J})$ while $|D|=\sum_{J\subset[m]}n_{J}|J|.$ When $i\in\Theta(J)$, we write $\widehat{L}_{\underline{g}}(i,J)$ instead of $L_{\underline{g}}(i,J)$ to emphasize that $L_{\underline{g}}(i,J)\in\widehat{D}.$
Remark 4.3.
The following “canonical” choice of $\Theta(J)\subset J$ is used in [GPTW16] and [PV16, PV19]: define $\Theta(J)$ to be the set of all $i\in J$ such that
  1. (1)
    The vertices $i$ and $\max(J)$ are in different path components of $|\mathcal{K}_{J}|;$
  2. (2)
    The vertex $i$ has the smallest number in its path component of $|\mathcal{K}_{J}|.$
Then the Panov–Veryovkin set of generators for the group $\operatorname{Cart}(\underline{G},\mathcal{K})$ is naturally indexed by the same set. Indeed, for $i\in J\subset[m]$ and ${\underline{g}}\in\prod_{j\in J}G_{j}^{*}$ define
$\Gamma_{\underline{g}}(i,J):=(g_{k_{1}},(g_{k_{2}},\dots(g_{k_{s}},(g_{\max(J)},g_{i}))\dots))\in\operatorname{Cart}(\underline{G},\mathcal{K}),$
where $J\setminus\{i,\max(J)\}=\{k_{1}<\dots<k_{s}\}$ and $(g,h):=g^{-1}h^{-1}gh$ is the group commutator. Then $\{\Gamma_{\underline{g}}(i,J):i\in\Theta(J),J\subset[m]\}$ is precisely the set of generators from [PV19, Theorem 5.2].
Panov–Veryovkin generators are important for the analogy between the group $\pi_{*}(\mathcal{R}_{\mathcal{K}})$ and the Hopf algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbb{Z})$ (see Subsection 6.4 below). However, our choice of generators is more suitable for computations.
Remark 4.4.
For right-angled Coxeter groups $\underline{G}^{\mathcal{K}}=(\underline{\mathbb{Z}_{2}})^{\mathcal{K}}=\mathrm{RC}_{\mathcal{K}}$ we omit the subscript ${\underline{g}}$ since then $|G_{j}^{*}|=1$. Li Cai proved ([Cai21], unpublished) that the set $\widehat{D}=\{\widehat{L}(i,J):J\subset[m],i\in\Theta(J)\}\subset\mathrm{RC}_{\mathcal{K}}^{\prime}$ generates the group $\mathrm{RC}_{\mathcal{K}}^{\prime}.$ His approach relies on a study of the Davis complex [Dav08, Chapter 7], a contractible cubical complex which is the universal covering of $\mathcal{R}_{\mathcal{K}}.$
4.2. Reduction to the distinguished elements
Lemma 4.5.
Let $J\subset[m]$, $\{i,j\}\in\mathcal{K}_{J}$ and ${\underline{g}}\in\prod_{j\in J}G_{j}^{*}.$ Then
(4.1) $L_{\underline{g}}(i,J)L_{\underline{g}}(j,J\setminus i)=L_{\underline{g}}(j,J)L_{\underline{g}}(i,J\setminus j)\in\operatorname{Cart}(\underline{G},\mathcal{K}).$
Proof.
Indeed, we have $L_{\underline{g}}(i,J)\cdot L_{\underline{g}}(j,J\setminus i)=\prod_{k\in J}g_{k}\cdot g_{i}^{-1}g_{j}^{-1}\cdot(\prod_{k\in J\setminus\{i,j\}}g_{k})^{-1},$ and the right hand side is symmetrical on $i,j$ since $g_{i}$ and $g_{j}$ commute. ∎
Also, $\prod_{j\in J}g_{j}=\prod_{j\in J\setminus\{\max(J)\}}g_{j}\cdot g_{\max(J)}$ and hence
(4.2) $L_{\underline{g}}(\max(J),J)=1\in\operatorname{Cart}(\underline{G},\mathcal{K}).$
Problem 4.6.
Apply the Reidemeister-Schreier algorithm [LS77, Section II.4] to the subgroup $\operatorname{Cart}(\underline{G},\mathcal{K})\subset\underline{G}^{\mathcal{K}}$. Is it true that one obtains the presentation of $\operatorname{Cart}(\underline{G},\mathcal{K})$ by the generators $\{L_{\underline{g}}(i,J):i\in J\subset[m],~{}{\underline{g}}\in\prod_{j\in J}G_{j}^{*}\}$ modulo the relations (4.1) and (4.2)? This larger presentation would be explicit and natural with respect to maps of both groups and simplicial complexes. Probably, Theorem 4.11 can be proved by applying Tietze transformations to this presentation. However, this approach seems tedious.
Proposition 4.7.
Every element $L_{\underline{g}}(i,J)\in D$ is equal in $\operatorname{Cart}(\underline{G},\mathcal{K})$ to a word $\operatorname{Red}(L_{\underline{g}}(i,J))\in F(\widehat{D})$ on the distinguished elements. This word can be chosen naturally with respect to inclusions $\mathcal{K}\hookrightarrow\mathcal{L}$ of full subcomplexes.
Proof.
We say that $|J|$ is the length of $L_{\underline{g}}(i,J)$, and argue by induction on length. The base cases $|J|\leq 2$: $L_{\underline{g}}(i,\{i\})=1;$ $L_{\underline{g}}(i,\{i,j\})=1$ if $\{i,j\}\in\mathcal{K};$ otherwise $L_{\underline{g}}(i,\{i,j\})=\widehat{L}_{\underline{g}}(i,\{i,j\})$ for $i<j$ and $L_{\underline{g}}(i,\{i,j\})=\widehat{L}_{\underline{g}}(j,\{i,j\})^{-1}$ for $i>j.$
At the inductive step, for an element $L_{\underline{g}}(i,J)\in D$ there are two cases:
  • $i$ and $\max(J)$ are in the same path component of $\mathcal{K}_{J}.$ Choose a path
    $(i=i_{0},i_{1},\dots,i_{k}=\max(J))$
    in $\mathcal{K}_{J},$ i.e. a sequence of edges $\{i_{t},i_{t+1}\}\in\mathcal{K}_{J}$. Applying the relation (4.1), we replace $L_{\underline{g}}(i_{t},J)$ with a product of $L_{\underline{g}}(i_{t+1},J)$ and generators of smaller length. On the last step we replace $L_{\underline{g}}(\max(J),J)$ with $1$ (using (4.2)). Therefore, $L_{\underline{g}}(i,J)$ is replaced with a word on elements of smaller length. By the inductive hypothesis, each of this elements can be expressed through the distinguished elements.
  • $i$ and $\max(J)$ are in different path components of $\mathcal{K}_{J}.$ Then there is exactly one vertex $i^{\prime}\in\Theta(J)$ such that $i$ and $i^{\prime}$ are in the same path component. Choose a path $(i=i_{0},\dots,i_{k}=i^{\prime})$ in $\mathcal{K}_{J}.$ Arguing as above, we obtain a word on $L_{\underline{g}}(i^{\prime},J)=\widehat{L}_{\underline{g}}(i^{\prime},J)$ and on elements of smaller length.
To obtain functoriality, choose each time the lexicographically minimal path from $i$ to $i_{k}.$ This choice is preserved by inclusions of simplicial complexes. ∎
Remark 4.8.
The algorithm of Proposition 4.7 does not depend on choice of ${\underline{g}}\in\prod_{j}G_{j}^{*}$ and on the group structure in $G_{1},\dots,G_{m}$. In more detail, if
$\operatorname{Red}(L_{\underline{g}}(i,J))=\widehat{L}_{\underline{g}}(i_{1},J_{1})^{\varepsilon_{1}}\cdot\dotso\cdot\widehat{L}_{\underline{g}}(i_{N},J_{N})^{\varepsilon_{N}}$
in $\operatorname{Cart}(\underline{G},\mathcal{K})$, then
$\operatorname{Red}(L_{\underline{h}}(i,J))=\widehat{L}_{\underline{h}}(i_{1},J_{1})^{\varepsilon_{1}}\cdot\dotso\cdot\widehat{L}_{\underline{h}}(i_{N},J_{N})^{\varepsilon_{N}}$
in $\operatorname{Cart}(\underline{H},\mathcal{K})$ for any $\underline{H}=(H_{1},\dots,H_{m})$ and ${\underline{h}}\in\prod_{j}H_{j}^{*}.$
4.3. Statement of the main theorem
Definition 4.9.
Let $W\in F(D)$. Express the letters of $W$ through the distinguished elements using the algorithm from Proposition 4.7. The resulting word will be denoted by $\operatorname{Red}(W)\in F(\widehat{D}).$
Definition 4.10.
Let $J\subset[m]$, ${\underline{g}}\in\prod_{j\in J}G_{j}^{*}$. Let $\lambda=(i_{1},\dots,i_{k+1}=i_{1})$ be a cycle in $\mathcal{K}_{J},$ i.e. a sequence of edges $\{i_{t},i_{t+1}\}\in\mathcal{K}_{J}$ for $t=1,\dots,k.$ Define the word
$R_{\underline{g}}(\lambda,J):=\prod_{t=1}^{k}L_{\underline{g}}(i_{t+1},J\setminus i_{t})\cdot L_{\underline{g}}(i_{t},J\setminus i_{t+1})^{-1}\in F(D).$
For example, if $J=\{4,5,7,8\}$ and $\lambda=(7,5,4,7),$ we have
$R_{\underline{g}}(\lambda,J)=L_{\underline{g}}(5,458)L_{\underline{g}}(7,478)^{-1}\cdot L_{\underline{g}}(4,478)L_{\underline{g}}(5,578)^{-1}\cdot L_{\underline{g}}(7,578)L_{\underline{g}}(4,458)^{-1}.$
[…]
Theorem 4.11.
Let $\underline{G}=(G_{1},\dots,G_{m})$ be a sequence of discrete groups and $\mathcal{K}$ be a flag simplicial complex on vertex set $[m].$ For each $J\subset[m],$ choose a set $\operatorname{Gen}(J)$ of cycles $\lambda=(i_{1},\dots,i_{k},i_{k+1}=i_{1})$ in $\mathcal{K}_{J}$ such that their images generate the group $\Pi_{1}(\mathcal{K}_{J})$.
Then the group $\operatorname{Cart}(\underline{G},\mathcal{K})$ is presented by $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot n_{J}$ generators
$\widehat{D}=\{\widehat{L}_{\underline{g}}(i,J):J\subset[m],~{}{\underline{g}}\in\prod_{j\in J}G_{j}^{*},~{}i\in\Theta(J)\}$
modulo $\sum_{J\subset[m]}|\operatorname{Gen}(J)|\cdot n_{J}$ relations
$\{\operatorname{Red}(R_{\underline{g}}(\lambda,J))=1:J\subset[m],{\underline{g}}\in\prod_{j\in J}G_{j}^{*},~{}\lambda\in\operatorname{Gen}(J)\}.$
The proof will be given in section 5.
Proof of Theorem 1.1.
By CW-approximation and Proposition 2.10(1), for each $J\subset[m]$ there exists a minimal generating set $\operatorname{Gen}(J)$ for the group $\Pi_{1}(\mathcal{K}_{J})$ represented by cycles, i.e. $|\operatorname{Gen}(J)|=\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$. The corresponding presentation from Theorem 4.11 meets the conditions of Theorem 1.1. ∎
Remark 4.12.
For ${\underline{g}}\in\prod_{j=1}^{m}G_{j}$, denote $\operatorname{supp}{\underline{g}}:=\{j\in[m]:~{}g_{j}\neq 1_{j}\}$. Then we can rewrite $\widehat{D}=\{\widehat{L}_{\underline{g}}(i,\operatorname{supp}{\underline{g}}):~{}{\underline{g}}\in\prod_{j=1}^{m}G_{j},~{}i\in\Theta(\operatorname{supp}{\underline{g}})\}$ and the set of relations as $\{\operatorname{Red}(R_{\underline{g}}(\lambda,\operatorname{supp}{\underline{g}}))=1:~{}{\underline{g}}\in\prod_{i=1}^{m}G_{j},~{}\lambda\in\operatorname{Gen}(\operatorname{supp}{\underline{g}})\}$.
4.4. Examples
Example 4.13 ($4$-cycle).
Let $\mathcal{K}$ be a four-cycle. Since $\mathcal{K}_{\{1,3\}}$ and $\mathcal{K}_{\{2,4\}}$ are the only disconnected full subcomplexes, $\Theta(\{1,3\})=\{1\},$ $\Theta(\{2,4\})=\{2\}$, all other $\Theta(J)$ being zero. We denote the corresponding distinguished elements $\widehat{L}_{\underline{g}}(1,\{1,3\})=g_{1}g_{3}\cdot g_{1}^{-1}\cdot g_{3}^{-1}$ and $\widehat{L}_{\underline{g}}(2,\{2,4\})=g_{2}g_{4}\cdot g_{2}^{-1}\cdot g_{4}^{-1}$ by $A_{g_{1},g_{3}}$ and $B_{g_{2},g_{4}}.$ It follows that the group $\operatorname{Cart}(\underline{G},\mathcal{K})=\operatorname{Ker}((G_{1}\ast G_{3})\times(G_{2}\ast G_{4})\to G_{1}\times G_{2}\times G_{3}\times G_{4})$ is generated by the set
$\widehat{D}=\widehat{A}\sqcup\widehat{B}=\{A_{g_{1},g_{3}}:g_{1}\in G_{1}^{*},g_{3}\in G_{3}^{*}\}\sqcup\{B_{g_{2},g_{4}}:g_{2}\in G_{2}^{*},g_{4}\in G_{4}^{*}\}$
modulo the following relations: for every ${\underline{g}}\in\prod_{j=1}^{4}G_{j}^{*},$ there is a relation
$\operatorname{Red}(R_{\underline{g}}(\lambda,\{1,2,3,4\}))=1,$ where $\lambda=(1,2,3,4,1).$ The left hand side is equal to
$\operatorname{Red}\Big{(}$ $L_{\underline{g}}(2,234)L_{\underline{g}}(1,134)^{-1}L_{\underline{g}}(3,134)L_{\underline{g}}(2,124)^{-1}$
$\cdot$ $L_{\underline{g}}(4,124)L_{\underline{g}}(3,123)^{-1}L_{\underline{g}}(1,123)L_{\underline{g}}(4,234)^{-1}\Big{)}=$
$\widehat{L}_{\underline{g}}(2,24)\cdot\widehat{L}_{\underline{g}}(1,13)^{-1}\cdot 1\cdot\widehat{L}_{\underline{g}}(2,24)^{-1}\cdot 1\cdot 1^{-1}\cdot\widehat{L}_{\underline{g}}(1,13)\cdot 1^{-1}=(B_{g_{2},g_{4}}^{-1},A_{g_{1},g_{3}}).$
Equivalently, $A_{g_{1},g_{3}}$ and $B_{g_{2},g_{4}}$ commute, hence $\operatorname{Cart}(\underline{G},\mathcal{K})\cong F(\widehat{A})\times F(\widehat{B})$ is a direct product of two free groups. In particular, $\operatorname{Cart}(\underline{\mathbb{Z}_{2}},\mathcal{K})=\mathrm{RC}^{\prime}_{\mathcal{K}}\cong\mathbb{Z}^{2}$ is generated by $A_{g_{1},g_{3}}$ and $B_{g_{2},g_{4}}.$
Example 4.14 ($m$-cycles).
Let $\mathcal{K}_{m}$ be an $m$-cycle, $m>4.$ By [BP15, Theorem 4.1.7], $\mathcal{R}_{\mathcal{K}}$ is a closed $(n+1)$-manifold whenever $\mathcal{K}$ is a triangulation of $n$-sphere. The $m$-cycle corresponds to the oriented surface $S_{g(m)}\cong\mathcal{R}_{\mathcal{K}_{m}}$ of genus $g(m)=1+(m-4)2^{m-3},$ hence $\mathrm{RC}^{\prime}_{\mathcal{K}_{m}}$ is the surface group $\pi_{1}(S_{g(m)}).$
Consider the presentation given by Theorem 4.11. The only full subcomplex with non-trivial $\Pi_{1}$ is the whole $\mathcal{K}_{m}.$ If the vertices are ordered cyclically, we have $\lambda=(1,2,\dots,m,1)$ and hence the following set of $\prod_{i=1}^{m}|G_{i}^{*}|$ defining relations for $\operatorname{Cart}(\underline{G},\mathcal{K}_{m}):$
$\Big{\{}\operatorname{Red}\Big{(}\prod_{t=1}^{m}L_{\underline{g}}(t+1,[m]\setminus\{t\})L_{\underline{g}}(t,[m]\setminus\{t+1\})^{-1}\Big{)}=1,~{}\forall{\underline{g}}\in\prod_{i=1}^{m}G_{i}^{*}\Big{\}}.$
In particular, $\mathrm{RC}^{\prime}_{\mathcal{K}_{m}}$ is defined by the single relation $R_{m}=1,$ where
$R_{m}:=\operatorname{Red}\Big{(}\prod_{t=1}^{m}L(t+1,[m]\setminus\{t\})L(t,[m]\setminus\{t+1\})^{-1}\Big{)}.$
For $m=5$ we have the relation
$\operatorname{Red}\Big{(}$ $L(2,2345)L(1,1345)^{-1}\cdot L(3,1345)L(2,1245)^{-1}$
$\cdot$ $L(4,1235)L(3,1235)^{-1}\cdot L(5,1235)L(4,1234)^{-1}\cdot L(1,1234)L(5,2345)^{-1}\Big{)}=$
$\widehat{L}(2,235)\widehat{L}(3,35)\widehat{L}(2,25)^{-1}\widehat{L}(2,245)\widehat{L}(3,35)^{-1}\widehat{L}(1,134)^{-1}$
$\cdot$ $\widehat{L}(3,135)\widehat{L}(1,14)\widehat{L}(2,245)^{-1}\widehat{L}(1,124)^{-1}\widehat{L}(2,25)\widehat{L}(3,135)^{-1}\widehat{L}(1,13)$
$\cdot$ $\widehat{L}(2,235)^{-1}\widehat{L}(1,13)^{-1}\widehat{L}(1,124)\widehat{L}(2,24)\widehat{L}(1,14)^{-1}\widehat{L}(1,134)\widehat{L}(2,24)^{-1}.$
[…]
Figure 1. The relation $R_{5}$. Notation: e.g. $1\underline{3}5^{-1}$ denotes $\widehat{L}(3,\{1,3,5\})^{-1}$.
Note that this element belongs to the commutator subgroup as predicted by Proposition 3.5. The presentation complex of this relation is obtained from the $20$-gon by gluing the edges as in Figure 1 while preserving the orientation. This complex is easily verified to be a surface of genus $5$.
The Python implementation of an algorithm which computes the relation $R_{m}$ (and, more generally, the relation $\operatorname{Red}(R_{\underline{g}}(\lambda,J))\in F(\widehat{D})$ for any given $\mathcal{K}$, $J$ and $\lambda$) is available on the author’s Github repository [Vyl24*].
The complexity of the reduction grows exponentially on $m,$ but the algorithm is relatively fast: we are able to calculate the relation $R_{m}$ for $m=20$ in less then $30$ minutes. This relation is a word of length $8388612$.
Remark 4.15.
Consider an arbitrary presentation of the surface group $\pi_{1}(S_{g})$ by $2g$ generators $x_{1},\dots,x_{2g}$ modulo one relation $R$. Every generator appears in $R$ at least twice, since $R$ belongs to the commutator subgroup of $F(x_{1},\dots,x_{2g})$ by Proposition 3.2. Therefore, $|R|\geq 4g$. In the standard presentation
$\pi_{1}(S_{g})\cong\langle a_{1},b_{1},\dots,a_{g},b_{g}\mid a_{1}b_{1}a_{1}^{-1}b_{1}^{-1}\cdot\dotso\cdot a_{g}b_{g}a_{g}^{-1}b_{g}^{-1}\rangle$
the relation satisfies $|R|=4g.$ Computer experiments show that, for the generators $\widehat{D}=\{\widehat{L}(i,J)\}$ of $\mathrm{RC}^{\prime}_{\mathcal{K}_{m}}\cong\pi_{1}(S_{g(m)}),$ the length of the relation $R_{m}$ given by Theorem 4.11 is also equal to $4g(m)$ for $m\leq 20$ (and, possibly, for all $m$). Probably, this can be explained by the geometric nature of the proof of Theorem 4.11: at the last step, we glue the disc ${\mathrm{cc}}(\mathcal{K})_{\underline{-1}}\simeq D^{2}$ to the punctured surface $(\mathcal{R}_{\mathcal{K}_{m}})^{\circ}:=\bigcup_{{\underline{g}}\neq\underline{-1}}{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ by the boundary, obtaining the closed surface $\mathcal{R}_{\mathcal{K}_{m}}$. We guess that the generators $\{\widehat{L}(i,J)\}$ of the free group $\pi_{1}((\mathcal{R}_{\mathcal{K}_{m}})^{\circ})$ can be represented by simple closed curves in $\mathcal{R}_{\mathcal{K}_{m}}$ and hence $\mathcal{R}_{\mathcal{K}_{m}}$ is homeomorphic to the presentation complex for this presentation.
Our computer experiments show that the single relation between the Panov–Veryovkin generators $\{\widehat{\Gamma}(i,J)\}$ has length much larger than $4g(m).$ For example, if $m=10$ then $4g(m)=3076,$ but the relation between the Panov–Veryovkin generators has length $63940$
5. Proof of Theorem 4.11
5.1. Outline of the proof
Consider the polyhedral product $B:=(\operatorname{cone}(\underline{G}),\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$. In Proposition 5.11 we give a geometric interpretation of the isomorphism $\pi_{1}(B)\cong\operatorname{Cart}(\underline{G},\mathcal{K})$, identifying certain loops in $B$ with elements of $\operatorname{Cart}(\underline{G},\mathcal{K})$. Then, following Li Cai [Cai21], we represent $B$ as the union of contractible subspaces ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ (see (5.1)), ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$.
For a subset $Q\subset\prod_{i=1}^{m}G_{i}$, denote $B_{Q}:=\bigcup_{{\underline{h}}\in Q}{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$. Then $B_{Q\sqcup\{{\underline{g}}\}}=B_{Q}\cup{\mathrm{cc}}(\mathcal{K})_{\underline{g}},$ where ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is contractible. In some cases the intersection $B_{Q}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is homotopy equivalent to $|\mathcal{K}_{J}|$ for $J=\operatorname{supp}{\underline{g}}\subset[m]$ (Proposition 5.9). Hence, by an application of the van Kampen theorem (Lemma 5.1), the group $\pi_{1}(B_{Q\sqcup\{{\underline{g}}\}})$ is obtained from $\pi_{1}(B_{Q})$ by imposing $\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})$ relations and then adding $\widetilde{b}_{0}(\mathcal{K}_{J})$ generators. By induction, in Theorem 5.14 we give presentations for fundamental groups of $\pi_{1}(B_{Q})$ whenever $Q\subset\prod_{i=1}^{m}G_{i}$ is an admissible subset (see Definition 5.6). A presentation of $\pi_{1}(B)$ is then obtained by passing to the limit (Lemma 5.21).
5.2. The geometric model for the classifying space
[…]
Figure 2. Polyhedral products $B=(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ as unions of ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$
By Proposition 2.5, the polyhedral product $(E\underline{G},\underline{G})^{\mathcal{K}}$ is the classifying space of the Cartesian subgroup $\operatorname{Cart}(\underline{G},\mathcal{K})$. Since $E\underline{G}$ is contractible, we have homotopy equivalences $(EG_{i},G_{i})\simeq(\operatorname{cone}G_{i},G_{i})$ and $(E\underline{G},\underline{G})^{\mathcal{K}}\simeq B$, where $B:=(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}=(\underline{X},\underline{A})^{\mathcal{K}}$ for $X_{i}:=\operatorname{cone}G_{i}$ and $A_{i}:=G_{i}$. Therefore, $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}=B(\operatorname{Cart}(\underline{G},\mathcal{K})).$
We denote the vertex of $\operatorname{cone}G_{i}$ by 00 and the segment between 00 and $g_{i}$ by $[0,g_{i}]$. Therefore $\operatorname{cone}G_{i}=\bigcup_{g_{i}\in G_{i}}[0,g_{i}]$. For each ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$, denote
(5.1) ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}:=([0,{\underline{g}}],\{{\underline{g}}\})^{\mathcal{K}}.$
More formally, ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}=(\underline{X},\underline{A})^{\mathcal{K}}$ for $X_{i}=[0,g_{i}],~{}A_{i}=\{g_{i}\}.$ It is clear that $B$ is glued out of $|\prod_{i=1}^{m}G_{i}|$ copies of ${\mathrm{cc}}(\mathcal{K})$,
$(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}=\bigcup\Big{\{}{\mathrm{cc}}(\mathcal{K})_{\underline{g}}:{\underline{g}}\in\prod_{i=1}^{m}G_{i}\Big{\}}.$
In particular, the real moment-angle complex $(\operatorname{cone}\mathbb{Z}_{2},\mathbb{Z}_{2})^{\mathcal{K}}=([-1,1],\{\pm 1\})^{\mathcal{K}}$ is glued out of $2^{m}$ copies of ${\mathrm{cc}}(\mathcal{K}).$
The polyhedral product $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ and its subspace ${\mathrm{cc}}(\mathcal{K})_{\underline{1}}$ is shown in Figure 2 for the case $\mathcal{K}=\bullet\bullet\!\!-\bullet$ and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{2},\mathbb{Z}_{2})$ (left) and $\underline{G}=(\mathbb{Z}_{2},\mathbb{Z}_{3},\mathbb{Z}_{2})$ (right).
Throughout the proof, we will study the fundamental groups of path connected subspaces $B_{Q}:=\bigcup_{{\underline{g}}\in Q}{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\subset(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ for certain $Q\subset\prod_{i=1}^{m}G_{i}$. Adding an element ${\underline{h}}$ to the set $Q$ corresponds to attaching a contractible space ${\mathrm{cc}}(\mathcal{K})_{\underline{h}}$ to the space $B_{Q}$ along a common CW-subcomplex. This affects the fundamental group as follows.
Lemma 5.1.
Let $Y$ be a CW complex and $X,C\subset Y$ be its subcomplexes such that $X$ is path connected, $C$ is simply connected and $Y=X\cup C$. Let $X\cap C=A=\bigsqcup_{\alpha\in PC}A_{\alpha}$ be the decomposition into path components. Then a presentation of $\pi_{1}(Y)$ can be obtained from an arbitrary presentation of $\pi_{1}(X)$ by adding $\operatorname{rank}\Pi_{1}(A)$ relations and then adding $\widetilde{b}_{0}(A)$ generators, as follows.
Let $x\in X$ be a basepoint. For each $\alpha\in PC,$ choose a basepoint $a_{\alpha}\in A_{\alpha}$, a path $t_{\alpha}$ in $X$ from $x$ to $a_{\alpha}$, and a set of loops $\{\lambda_{i,\alpha}:i\in I_{\alpha}\}$ in $A_{\alpha}$ that generate the group $\pi_{1}(A_{\alpha},a_{\alpha}).$
Choose a distinguished element $0\in PC.$ For each $\alpha\in PC\setminus\{0\},$ choose a path $s_{\alpha}$ in $C$ from $a_{0}$ to $a_{\alpha}.$
Then $\pi_{1}(Y,x)$ is the free product of the quotient group
$\pi_{1}(X,x)/(t_{\alpha}\cdot\lambda_{i,\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC,~{}i\in I_{\alpha})$
with the free group generated by the loops $\{t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\}\}.$
[…]
Figure 3. The space $Y=X\cup_{A}C$ from Lemma 5.1.
Proof.
Without loss of generality, $x$ and $a_{\alpha}$ are 0-cells and $s_{\alpha}$ and $t_{\alpha}$ are distinct 1-cells of $Y.$ (If this is not the case, we first use cellular approximation theorem to ensure that images of $s_{\alpha}$ and $t_{\alpha}$ are in $\operatorname{sk}_{1}Y,$ and then attach a 1-cell and a 2-cell to each path without changing homotopy types, compare with the proof of [Hat02, Proposition 1.26].) The closures of these cells are CW-subcomplexes $\overline{t_{\alpha}}=\{x,a_{\alpha}\}\cup t_{\alpha}\subset X$ and $\overline{s_{\alpha}}=\{a_{0},a_{\alpha}\}\cup s_{\alpha}\subset C.$
Now denote $T:=\overline{t_{0}}\cup\bigcup_{\alpha\in PC\setminus\{0\}}\overline{s_{\alpha}}.$ We have the following pushout diagram of inclusion of connected CW-complexes with the basepoint $x$:
${A\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${X\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C\cup T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y.}$
Now we describe their fundamental groups and maps between them.
  • Clearly, $(A\cup T)/T\cong\bigvee_{\alpha\in PC}A_{\alpha}.$ Since $T$ is contractible, we have
[…]
    The isomorphism $\Pi_{1}(A)\overset{\cong}{\longrightarrow}\pi_{1}(A\cup T,x)$ is described as follows: if $\lambda$ is a loop in $A_{\alpha}$ with the basepoint $a_{\alpha},$ then $t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1}$ is the corresponding loop in $A\cup T$ with the basepoint $x.$
  • The space $C\cup T$ is obtained from a simply connected space $C$ by attaching a segment $[a_{0},x]$ at the point $a_{0}\in C$, hence $\pi_{1}(C\cup T,x)$ is trivial.
  • The space $X\cup T$ is obtained from $X$ by attaching 1-cells $\{s_{\alpha}:\alpha\in PC\setminus\{0\}\}$ connecting the points $a_{0}$ and $a_{\alpha}.$ Clearly, $X\cup T\simeq(X\cup T)/T\simeq X\vee\bigvee_{\alpha\in PC\setminus\{0\}}S^{1}.$ Hence $\pi_{1}(X\cup T)$ is the free product of $\pi_{1}(X)$ with the free group on $\widetilde{b}_{0}(A)$ generators,
    $\pi_{1}(X\cup T,x)\cong\pi_{1}(X,x)\ast F(\{t_{0}\cdot s_{\alpha}\cdot t_{\alpha}^{-1}:\alpha\in PC\setminus\{0\}\}).$
Applying the van Kampen theorem to the diagram above, we obtain the following pushout in the category of groups:
[…]
where $i_{*}$ is the free product of maps $\pi_{1}(A,a_{\alpha})\to\pi_{1}(X),$ $\lambda\mapsto t_{\alpha}\cdot\lambda\cdot t_{\alpha}^{-1}.$
5.3. Subspaces of ${\mathrm{cc}}(\mathcal{K})$
The barycentric subdivision of a simplicial complex $\mathcal{K}$ is the following simplicial complex $\mathcal{K}^{\prime}$ on the vertex set $\mathcal{K}\setminus\{\varnothing\}$,
$\mathcal{K}^{\prime}:=\Big{\{}\{I_{1},\dots,I_{r}\}:~{}r\geq 0;~{}I_{j}\in\mathcal{K},~{}j=1,\dots,r;~{}I_{1}\subsetneq I_{2}\subsetneq\dots\subsetneq I_{r};~{}I_{1}\neq\varnothing\Big{\}}.$
Hence the faces of $\mathcal{K}^{\prime}$ correspond to the chains $(I_{1}\subsetneq\dots\subsetneq I_{r})$ of non-empty simplices in $\mathcal{K}$. The vertices of $\mathcal{K}^{\prime}$ correspond to the non-empty simplices $I\in\mathcal{K};$ we denote them by $(I)\in\mathcal{K}^{\prime}$ and call the barycenters. There is a natural PL homeomorphism $\beta:|\mathcal{K}^{\prime}|\overset{\cong}{\longrightarrow}|\mathcal{K}|$ which maps $(I)\in\mathcal{K}^{\prime}$ to the geometric barycenter $b_{I}$ of the geometric simplex $|I|=\operatorname{conv}(e_{i}:i\in I)\subset|\mathcal{K}|.$
The simplicial cone over $\mathcal{K}^{\prime}$ is naturally identified with the simplicial complex
$\operatorname{cone}(\mathcal{K}^{\prime}):=\{(I_{1}\subsetneq\dots\subsetneq I_{r}):r\geq 0;~{}I_{j}\in\mathcal{K},~{}j=1,\dots,r\}$
on the vertex set $\mathcal{K}.$ Note that ()($\varnothing$)( ∅ ) is a vertex of $\operatorname{cone}(\mathcal{K}^{\prime})$, and its geometric realisation is the apex $b_{\varnothing}$ of the geometric cone $|\operatorname{cone}(\mathcal{K}^{\prime})|$ over $|\mathcal{K}^{\prime}|$ (“the barycenter of the empty simplex”).
Now let $e_{1},\dots,e_{m}$ be the standard basis of $\mathbb{R}^{m},$ and denote $e_{I}:=\sum_{i\in I}e_{i}\in[0,1]^{m}$ for $I\subset[m].$
Construction 5.2 ([BP15, §2.9]).
Let $\mathcal{K}$ be a simplicial complex. Consider the PL embedding $\iota_{\mathcal{K}}:|\operatorname{cone}(\mathcal{K}^{\prime})|\to[0,1]^{m}$ which is defined on vertices as $b_{I}\mapsto e_{[m]\setminus I}$ and extends linearly onto simplices of $|\operatorname{cone}(\mathcal{K}^{\prime})|$.
Lemma 5.3 (see [BP15, Proposition 2.9.12]).
The map $\iota_{\mathcal{K}}$ induces a homeomorphism $|\operatorname{cone}(\mathcal{K}^{\prime})|\overset{\cong}{\longrightarrow}{\mathrm{cc}}(\mathcal{K})\subset[0,1]^{m}$, where ${\mathrm{cc}}(\mathcal{K})=([0,1],\{1\})^{\mathcal{K}}.$
Recall that, for a simplex $A$ of a simplicial complex $\mathcal{L},$ the subcomplex
$\operatorname{st}_{\mathcal{L}}A:=\{I\in\mathcal{K}:I\cup A\in\mathcal{L}\}$
of $\mathcal{L}$ is called the star of $J$. In the next two lemmas, we consider $\mathcal{L}=\mathcal{K}^{\prime}$ and vertices ($(\{j\})$ of $\mathcal{K}^{\prime}$ that correspond to vertices $j$ of $\mathcal{K}$.
Lemma 5.4.
For each $j\in[m]$, we have $\iota_{\mathcal{K}}^{-1}(\{x_{j}=0\})=|\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\})|.$
Proof.
Since $\iota_{\mathcal{K}}:|\operatorname{cone}(\mathcal{K}^{\prime})|\to[0,1]^{m}$ is linear on simplices, the preimage of $H_{j}:=\{x\in[0,1]^{m}:x_{j}=0\}$ under $\iota_{\mathcal{K}}$ is (the geometric realisation of) the full subcomplex of $\operatorname{cone}(\mathcal{K}^{\prime})$ spanned by the vertices that are mapped into $H_{j}$.
A vertex $b_{I}$ of $\operatorname{cone}(\mathcal{K}^{\prime})$ is mapped to $e_{[m]\setminus I},$ hence $\iota_{\mathcal{K}}(b_{I})\in H_{j}$ if and only if $j\in I$. It follows that $\iota_{\mathcal{K}}^{-1}(H_{j})$ is the geometric realisation of the simplicial complex $\{(I_{1}\subsetneq\dots\subsetneq I_{r})\in\operatorname{cone}(\mathcal{K}^{\prime}):j\in I_{1}\},$ which clearly coincides with $\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\}).$
Lemma 5.5.
For $J\subset[m],$ denote $\mathcal{N}(J):=\bigcup_{j\in J}\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\})\subset\mathcal{K}^{\prime}.$ Then
${\mathrm{cc}}(\mathcal{K})\cap\bigcup_{j\in J}\{x_{j}=0\}=\iota_{\mathcal{K}}(|\mathcal{N}(J)|)\simeq|\mathcal{K}_{J}|.$
Proof.
By the previous lemma, we have ${\mathrm{cc}}(\mathcal{K})\cap\{x_{j}=0\}=\iota_{\mathcal{K}}(|\operatorname{st}_{\mathcal{K}^{\prime}}(\{j\})|)$, so we obtain the first statement by taking the union over $j\in J$.
Since $\iota_{\mathcal{K}}:|\operatorname{cone}(\mathcal{K}^{\prime})|\to{\mathrm{cc}}(\mathcal{K})$ is a homeomorphism, now it is sufficient to prove that $|\mathcal{N}|$ and $|\mathcal{K}_{J}|$ are homotopy equivalent. Note that $\beta:|\mathcal{K}|\to|\mathcal{K}^{\prime}|$ is a PL homeomorphism, and $\beta(|\mathcal{K}_{J}|)\subset|\mathcal{N}|$.
Consider the full subcomplex of $\mathcal{K}^{\prime}$ on the vertex set $\{(I):I\in\mathcal{K},I\cap J\neq\varnothing\}$. Since $I\cap J\neq\varnothing$ if and only if $\{j\}\subsetneq I$ for some $j\in J$, this full subcomplex coincides with $\mathcal{N}$.
Consider the full subcomplex of $\mathcal{K}^{\prime}$ on the vertex set $\{(I):I\in\mathcal{K},I\subset J\}$. This vertex set is equal to $\mathcal{K}_{J}\setminus\{\varnothing\},$ hence the geometric realisation of this full subcomplex coincides with $\beta(|\mathcal{K}_{J}|)$.
Hence we can define a strict deformation retraction of $|\mathcal{N}|$ onto $\beta(|\mathcal{K}_{J}|)$ as follows: for every $I\in\mathcal{K}$ such that $I\cap J\neq\varnothing,$ we move $b_{I}$ linearly to the point $b_{I\cap J}$ along the segment $|(I\cap J\subsetneq I)|\subset|\mathcal{N}|;$ we extend this homotopy linearly onto geometric simplices of $|\mathcal{N}|$. This homotopy is well defined: if $(I_{1}\subsetneq\dots\subsetneq I_{r})\in\mathcal{N}$, then the geometric simplex $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}})\subset|\mathcal{N}|$ is linearly deformed onto the geometric simplex $\operatorname{conv}(b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})$; the image of the homotopy lies in $\operatorname{conv}(b_{I_{1}},\dots,b_{I_{r}},b_{I_{1}\cap J},\dots,b_{I_{r}\cap J})\subset\beta(|I_{r}|)\subset|\mathcal{N}|.$
5.4. Admissible subsets
For ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$, denote $\operatorname{supp}{\underline{g}}:=\{i\in[m]:g_{i}\neq 1_{i}\}.$
For ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$ and $I\subset[m]$, denote
${\underline{g}}(I):=(x_{1},\dots,x_{m})\in\prod_{i=1}^{m}G_{i},~{}x_{i}=\begin{cases}g_{i},&i\in I;\\ 1_{i},&i\notin I.\end{cases}$
Definition 5.6.
A subset $Q\subset\prod_{i=1}^{m}G_{i}$ is admissible if $Q$ is finite, non-empty, and satisfies ${\underline{g}}(I)\in Q$ whenever ${\underline{g}}\in Q$ and $I\subset[m]$. If $Q$ is admissible, an element ${\underline{g}}\in Q$ is maximal if we have ${\underline{g}}\neq{\underline{h}}(I)$ for all ${\underline{h}}\in Q\setminus\{{\underline{g}}\}$.
Consider the following partial ordering on $\prod_{i=1}^{m}G_{i}:$ ${\underline{g}}\leqslant{\underline{h}}$ if and only if ${\underline{g}}={\underline{h}}(I)$ for some $I\subset[m]$. Clearly, an element ${\underline{g}}\in Q$ is maximal if and only if ${\underline{g}}$ is a maximal element with respect to $\leqslant$. Also, $Q$ is admissible if and only if $Q$ is finite and is a lower set (i.e. ${\underline{g}}\in Q$ whenever ${\underline{g}}\leqslant{\underline{h}}\in Q$). Hence the following properties are clear.
Lemma 5.7.
  1. (1)
    A finite union of admissible subsets is admissible.
  2. (2)
    For any ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$, the subset $\{{\underline{g}}(I):I\subset[m]\}$ is admissible.
  3. (3)
    If $Q$ is admissible and ${\underline{g}}\in Q$ is maximal, then $Q\setminus\{{\underline{g}}\}$ is admissible.∎
For an admissible subset $Q\subset\prod_{i=1}^{m}G_{i}$, we denote $B_{Q}:=\bigcup_{{\underline{g}}\in Q}{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
Denote the composition of homeomorphisms $\iota_{\mathcal{K}}:|\operatorname{cone}(\mathcal{K}^{\prime})|\to{\mathrm{cc}}(\mathcal{K})$ and ${\mathrm{cc}}(\mathcal{K})\cong{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ by $\iota_{\underline{g}}$. In our notation, for $I\in\mathcal{K}$ we have
$\iota_{\underline{g}}(b_{I})=(x_{1},\dots,x_{m}),\quad x_{i}=\begin{cases}g_{i},~{}i\notin I;\\ 0,~{}i\in I.\end{cases}$
Lemma 5.8.
Let $x\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$ for some ${\underline{g}},{\underline{h}}\in\prod_{i=1}^{m}G_{i}$, and suppose that $x_{j}\neq 0$ for all $j\in\operatorname{supp}{\underline{g}}$. Then ${\underline{g}}={\underline{h}}(S)$ for $S=\{i\in[m]:x_{i}\neq 0\}$.
Proof.
If $i\in S$, then $x_{i}\neq 0$. Since $x\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$, we have $x_{i}\in[0,g_{i}]\cap[0,h_{i}]$; it follows that $g_{i}=h_{i}$.
If $j\notin S$, then $x_{j}=0$. We have $j\notin\operatorname{supp}{\underline{g}}$ by the assumption, hence $g_{j}=1_{j}$. ∎
Proposition 5.9.
Let ${\underline{g}}$ be a maximal element of an admissible subset $Q\subset\prod_{i=1}^{m}G_{i}$. Then $B_{Q\setminus\{{\underline{g}}\}}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=\iota_{\underline{g}}(|\mathcal{N}(\operatorname{supp}{\underline{g}})|)$. In particular, this space is homotopy equivalent to $\mathcal{K}_{\operatorname{supp}{\underline{g}}}$ by Lemma 5.5.
Proof.
Denote $J:=\operatorname{supp}{\underline{g}}$ and $X=B_{Q\setminus\{{\underline{g}}\}}.$ In the view of Lemma 5.5, it is sufficient to prove that $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=\bigcup_{j\in J}\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}.$
First let $j\in J$. Since $Q$ is admissible, we have ${\underline{g}}(J\setminus j)\in Q$, hence ${\mathrm{cc}}(\mathcal{K})_{{\underline{g}}(J\setminus j)}\subset X$. It is straightforward to show that ${\mathrm{cc}}(\mathcal{K})_{{\underline{g}}(J\setminus j)}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$. Thus $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\supseteq\bigcup_{j\in J}\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}.$
On the other hand, let $x=(x_{1},\dots,x_{m})\in X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}},$ and suppose that $x_{j}\neq 0$ for all $j\in J$. We have $x\in{\mathrm{cc}}(\mathcal{K})_{\underline{h}}$ for some ${\underline{h}}\in Q\setminus\{{\underline{g}}\}$. By Lemma 5.8, ${\underline{g}}={\underline{h}}(S)$ for some $S\subset[m]$. This contradicts the maximality of ${\underline{g}}$. Hence $x_{j}=0$ for some $j\in J$. We proved that $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\subseteq\bigcup_{j\in J}\{x_{j}=0\}\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}.$
5.5. Loops defined by words
Here we give a geometric interpretation of the isomorphism $\pi_{1}((\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}},\underline{1})\cong\operatorname{Cart}(\underline{G},\mathcal{K})$ from Proposition 2.5.
For each $k_{i},g_{i}\in G_{i}$, consider a standard path in the topological space $\operatorname{cone}G_{i}$ from the point $k_{i}$ to the point $k_{i}g_{i}$ along two straight segments, $k_{i}\leadsto 0\leadsto k_{i}g_{i}$.
Definition 5.10.
Given a sequence of elements $g_{i_{1}},g_{i_{2}},\dots,g_{i_{N}}$ ($i_{t}\in[m]$, $g_{i_{t}}\in G_{i_{t}}$), we define the following path in the space $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ which starts at the point $\underline{1}=(1_{1},\dots,1_{m})$:
$(1_{1},\dots,1_{i_{1}},\dots,1_{i_{2}},\dots)\leadsto(1_{1},\dots,g_{i_{1}},\dots,1_{i_{2}},\dots)\leadsto(1_{1},\dots,g_{i_{1}},\dots,g_{i_{2}},\dots)\leadsto\dots.$
Here at each step $t=1,\dots,N$ we change the $i_{t}$-th coordinate from $k_{i_{t}}$ to $k_{i_{t}}g_{i_{t}}$ by going along a standard path, while all the other coordinates are fixed. Hence the path ends at the point $(h_{1},\dots,h_{m})\in(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, where $h_{i}=\prod_{t:i_{t}=i}g_{i_{t}}\in G_{i}$.
Proposition 5.11.
Let $g_{i_{1}},\dots,g_{i_{N}}$ be a sequence of letters such that $\prod_{t:i_{t}=i}g_{i_{t}}=1_{i}$ for each $i\in[m]$. Let $\lambda:[0,1]\to(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ be the corresponding path from $\underline{1}$ to $\underline{1}$ given in Definition 5.10. Then the isomorphism $\pi_{1}((\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}},\underline{1})\to\operatorname{Cart}(\underline{G},\mathcal{K})$ from Proposition 2.5 maps $[\lambda]$ to $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}.$
Proof.
The element $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in G_{1}\ast\dots\ast G_{m}\cong\pi_{1}(BG_{1}\vee\dots\vee BG_{m})$ is represented by the concatenation of loops which correspond to elements $g_{i}\in G_{i}\cong\pi_{1}(BG_{i})$. Since the isomorphism $\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})\cong\underline{G}^{\mathcal{K}}$ from Proposition 2.5 is natural with respect to maps of simplicial complexes, the word $g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}\cong\pi_{1}((B\underline{G},\ast)^{\mathcal{K}})$ is represented by the same loop. On the other hand, this loop is the composition of $\lambda:[0,1]\to(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ with the homotopy equivalence $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}\simeq(E\underline{G},\underline{G})^{\mathcal{K}}$ and the natural projection $(E\underline{G},\underline{G})^{\mathcal{K}}\to(B\underline{G},\ast)^{\mathcal{K}}$. This projection is injective on fundamental groups, so the claim follows. ∎
Remark 5.12.
In fact, the construction above provides a bijective correspondence
$\{x\in\underline{G}^{\mathcal{K}}:\pi(x)={\underline{h}}\}\overset{\cong}{\longrightarrow}\{\text{paths in }(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}\text{ from }\underline{1}\text{ to }{\underline{h}}\}/\sim,~{}\forall{\underline{h}}\in\prod_{i=1}^{m}G_{i},$
where $\pi:\underline{G}^{\mathcal{K}}\to\prod_{i=1}^{m}G_{i}$ is the natural projection and $\sim$ is the homotopy modulo endpoints. Different choices of decomposition $x=g_{i_{1}}\cdot\dotso\cdot g_{i_{N}}\in\underline{G}^{\mathcal{K}}$ correspond to homotopic paths: for example, each edge $\{i,j\}\in\mathcal{K}$ gives rise to the relation $g_{i}g_{j}=g_{j}g_{i}$ in $\underline{G}^{\mathcal{K}}$ (hence to uncertainty in the decomposition) and to disjoint union of squares in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ (hence to homotopies between corresponding paths). We do not prove this fact since we do not use it in full generality.
Definition 5.13.
To each standard element $L_{\underline{g}}(i,J)\in D$ we assign the loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ given in Definition 5.10 using the sequence from Definition 4.1 (e.g. $L_{\underline{g}}(3,\{1,3,4\})$ defines the same path as the sequence $g_{1}g_{3}g_{4}g_{3}^{-1}g_{4}^{-1}g_{1}^{-1}$). Hence each word $w=L_{\underline{g}}(i_{1},J_{1})^{\pm 1}\cdot\dots\cdot L_{\underline{g}}(i_{N},J_{N})^{\pm 1}$ on such elements defines a loop in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$. By the proposition above, this loop corresponds to $w\in\operatorname{Cart}(\underline{G},\mathcal{K})$ under the isomorphism $\pi_{1}(B,\underline{1})\cong\operatorname{Cart}(\underline{G},\mathcal{K})$.
For a subspace $X\subset(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$, we say that the word $w$ defines the path in $X$ if the corresponding path in $(\operatorname{cone}\underline{G},\underline{G})^{\mathcal{K}}$ lies in $X$. Hence if $L_{\underline{g}}(i,J)$ defines a path in $X$, then it can be considered as an element in $\pi_{1}(X,\underline{1})$.
5.6. Fundamental groups of subspaces corresponding to admissible subsets
Theorem 5.14.
Let $Q\subset\prod_{i=1}^{m}G_{i}$ be an admissible subset. For each ${\underline{g}}\in Q$, choose a set of cycles $\operatorname{Gen}({\underline{g}})$ in $\mathcal{K}_{\operatorname{supp}{\underline{g}}}$ that generate the group $\Pi_{1}(\mathcal{K}_{\operatorname{supp}{\underline{g}}}).$
Then the group $\pi_{1}(B_{Q},\underline{1})$ is presented by $\sum_{{\underline{g}}\in Q}\widetilde{b}_{0}(\mathcal{K}_{\operatorname{supp}{\underline{g}}})$ generators, which correspond to words
$\{\widehat{L}_{\underline{g}}(i,\operatorname{supp}{\underline{g}}):~{}{\underline{g}}\in Q,~{}i\in\Theta(\operatorname{supp}{\underline{g}})\},$
modulo $\sum_{{\underline{g}}\in Q}|\operatorname{Gen}({\underline{g}})|$ relations
$\{\operatorname{Red}(R_{\underline{g}}(\lambda,\operatorname{supp}{\underline{g}}))=1:{\underline{g}}\in Q,~{}\lambda\in\operatorname{Gen}({\underline{g}})\}.$
($R_{\underline{g}}(\lambda,\operatorname{supp}{\underline{g}})\in F(D)$ and $\operatorname{Red}:F(D)\to F(\widehat{D})$ were defined in Subsection 4.3.)
Remark 5.15.
In particular, we claim that the word $\operatorname{Red}(R_{\underline{g}}(\lambda,\operatorname{supp}{\underline{g}}))\in F(\widehat{D})$ depends only on the generators $\{\widehat{L}_{\underline{h}}(i,\operatorname{supp}{\underline{h}}):{\underline{h}}\in Q,i\in\Theta(\operatorname{supp}{\underline{h}})\}\subset\widehat{D}.$
Proof of Theorem 5.14.
Induction on $|Q|$. The base case is $Q=\{\underline{1}\}$, when the presentation is empty and $B_{Q}={\mathrm{cc}}(\mathcal{K})_{\underline{1}}$ is contractible.
On the induction step, we choose a maximal element ${\underline{g}}\in Q$, and denote $X:=B_{Q\setminus\{{\underline{g}}\}}$, $J:=\operatorname{supp}{\underline{g}}$. Hence $B_{Q}=X\cup{\mathrm{cc}}(\mathcal{K})_{\underline{g}},$ and $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\simeq|\mathcal{K}_{J}|$ by Proposition 5.9. It is sufficient to prove that a presentation of $\pi_{1}(B_{Q},\underline{1})$ can be obtained from a presentation of $\pi_{1}(X,\underline{1})$ by adding the $\widetilde{b}_{0}(\mathcal{K}_{J})$ generators $\{\widehat{L}_{\underline{g}}(i,J):i\in\Theta(J)\}$ and the $|\operatorname{Gen}({\underline{g}})|$ relations $\{\operatorname{Red}(R_{\underline{g}}(\lambda,J)):\lambda\in\operatorname{Gen}({\underline{g}})\}$.
We first define some paths in $B_{Q}$ and show that, up to homotopy, their concatenations are defined by words on elements from $\widehat{D}\subset\operatorname{Cart}(\underline{G},\mathcal{K})$.
  • For each $I\subset[m]$ consider the point $v_{I}:={\underline{g}}([m]\setminus I)\in B_{Q}$ with the coordinates $(v_{I})_{j}=\begin{cases}1_{j},&j\in I\\ g_{j},&j\notin I\end{cases}$ and the path
    $p(I):(1_{1},\dots,1_{m})\leadsto((v_{I})_{1},1_{2},\dots,1_{m})\leadsto\dots\leadsto((v_{I})_{1},\dots,(v_{I})_{m})$
    in $B_{Q}$ from $\underline{1}$ to $v_{I}$. (Both $v_{I}$ and $p(I)$ are in $B_{Q}$, since $Q$ is admissible and ${\underline{g}}\in Q$. Moreover, for $I\neq\varnothing$ both $v_{I}$ and $p(I)$ are in $X$.)
  • For each $i\in[m]$ consider the point $a_{i}\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ with the coordinates $(a_{i})_{k}:=\begin{cases}0,&k=i\\ g_{k},&k\neq i\end{cases}$ and the path
    $t(i):\underbrace{(1_{1},\dots,1_{m})}_{=\underline{1}}\overset{p(\{i\})}{\leadsto}\underbrace{(g_{1},\dots,1_{i},\dots,g_{m})}_{=v_{\{i\}}}\leadsto\underbrace{(g_{1},\dots,0_{i},\dots,g_{m})}_{=a_{i}}$
    in $X$ from $\underline{1}$ to $a_{i}$. (Here the second part is the standard path along the segment $[1_{i},0_{i}]$.)
  • For each $i\in J$ consider the path $s_{i}:a_{\max(J)}\leadsto v_{\varnothing}\leadsto a_{i}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ from $a_{\max(J)}$ to $a_{i}$. (Here $s_{i}$ the concatenation of standard paths along the segments $[0_{\max(J)},g_{\max(J)}]$ and $[g_{i},0_{i}]$).
Lemma 5.16.
For $i\in J\subset[m]$, denote $m^{\prime}=\max(J).$ Then the loop $t(m^{\prime})\cdot s_{i}\cdot t(i)^{-1}:\underline{1}\leadsto a_{m^{\prime}}\leadsto a_{i}\leadsto\underline{1}$ defines the element in $\pi_{1}(B_{Q},\underline{1})$ represented by the word $L_{\underline{g}}(i,J)$.
Proof.
By definition, this loop is the following composition of paths:
$\underline{1}=(1_{1},\dots,1_{i},\dots,1_{m^{\prime}},\dots,1_{m})\overset{p(\{m^{\prime}\})}{\leadsto}(g_{1},\dots,g_{i},\dots,1_{m^{\prime}},\dots,g_{m})\leadsto$
$(g_{1},\dots,g_{i},\dots,0_{m^{\prime}},\dots,g_{m})\leadsto(g_{1},\dots,g_{i},\dots,g_{m^{\prime}},\dots,g_{m})\leadsto$
$(g_{1},\dots,0_{i},\dots,g_{m^{\prime}},\dots,g_{m})\leadsto(g_{1},\dots,1_{i},\dots,g_{m^{\prime}},\dots,g_{m})\leadsto\dots\leadsto\underline{1}.$
Since $g_{j}=1_{j}$ for $j>m^{\prime}$, this path is defined by the sequence of letters $g_{1}g_{2}\dots g_{m^{\prime}}\cdot g_{i}^{-1}\cdot g_{m^{\prime}}^{-1}\dots g_{i+1}^{-1}g_{i-1}^{-1}\dots g_{1}^{-1}=L_{\underline{g}}(i,J).$
$t(i)$$t(k)$$\underline{1}$$v_{\{i\}}$$v_{\{m^{\prime}\}}$$v_{\varnothing}$$a_{i}$$a_{m^{\prime}}$$s(i)$
Figure 4. Proof of Lemma 5.16
  • For each edge $\{i,j\}\in\mathcal{K}_{J}$, consider the point $u_{ij}\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ with the coordinates $(u_{ij})_{k}:=\begin{cases}g_{k},&k\neq i,j\\ 0,&k\in\{i,j\}\end{cases}$ and the path $\ell(i,j):a_{i}\leadsto u_{ij}\leadsto a_{j}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
Lemma 5.17.
Let $\{i,j\}\in\mathcal{K}_{J}$. Then the loop $t(i)\cdot\ell(i,j)\cdot t(j)^{-1}:\underline{1}\leadsto a_{i}\leadsto a_{j}\leadsto\underline{1}$ defines the element in $\pi_{1}(X,\underline{1})$ represented by the word $L_{\underline{g}}(j,J\setminus i)\cdot L_{\underline{g}}(i,J\setminus j)^{-1}$.
Proof.
By definition, $t(i)\cdot\ell(i,j)\cdot t(j)^{-1}$ is the concatenation $\underline{1}\overset{p(\{i\})}{\leadsto}v_{\{i\}}\leadsto a_{i}\leadsto u_{ij}\leadsto a_{j}\leadsto v_{\{j\}}\overset{p(\{j\})^{-1}}{\leadsto}\underline{1}.$ The paths $v_{\{i\}}\leadsto a_{i}\leadsto u_{ij}\leadsto a_{j}\leadsto v_{\{j\}}$ and $v_{\{i\}}\leadsto v_{\{i,j\}}\leadsto v_{\{j\}}$ are homotopic in $X$, since there are three squares
$\prod_{k\neq i,j}\{g_{k}\}\times([0_{i},1_{i}]\times[0_{j},1_{j}]\cup[0_{i},1_{i}]\times[0_{j},g_{j}]\cup[0_{i},g_{j}]\times[0_{j},1_{j}])\subset X.$
By adding and removing the path $p(\{i,j\})^{-1}\cdot p(\{i,j\})$ from $v_{\{i,j\}}$ to $\underline{1}$ and back, we see that our loop is homotopic in $X$ to the loop
$\underline{1}\overset{p(\{i\})}{\leadsto}v_{\{i\}}\leadsto v_{\{i,j\}}\overset{p(\{i,j\})^{-1}}{\leadsto}\underline{1}\overset{p(\{i,j\})}{\leadsto}v_{\{i,j\}}\leadsto v_{\{j\}}\overset{p(\{j\})}{\leadsto}\underline{1}$
which is represented by the word
$\prod_{k\neq i}g_{i}\cdot g_{j}^{-1}\cdot(\prod_{k\neq i,j}g_{k})^{-1}\cdot\prod_{k\neq i,j}g_{k}\cdot g_{i}\cdot(\prod_{k\neq j}g_{j})^{-1}=L_{\underline{g}}(J\setminus i,j)\cdot L_{\underline{g}}(J\setminus j,i)^{-1}.\qed$
$\underline{1}$$v_{\{i,j\}}$$v_{\{i\}}$$v_{\{j\}}$$v_{\varnothing}$$u_{i,j}$$a_{i}$$a_{j}$$(1_{i},1_{j})$$(g_{i},g_{j})$$(0_{i},0_{j})$$p(\{i\})$$p^{-1}(\{j\})$
$\underline{1}$$v_{\{i,j\}}$$v_{\{i\}}$$v_{\{j\}}$$v_{\varnothing}$$u_{i,j}$$a_{i}$$a_{j}$$g_{j}^{-1}$$p^{-1}(\{i,j\})$
Figure 5. Proof of Lemma 5.17
  • For each simplicial cycle $\lambda=(i_{1},\dots,i_{k},i_{k+1}=i_{1})$ in $\mathcal{K}_{J}$, consider the path $\ell(\lambda):a_{i_{1}}\overset{\ell(i_{1},i_{2})}{\leadsto}a_{i_{2}}\leadsto\dots\leadsto a_{i_{k}}\overset{\ell(i_{k},i_{1})}{\leadsto}a_{i_{1}}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
Lemma 5.18.
Let $\lambda=(i_{1},\dots,i_{k},i_{k+1}=i_{1})$ be a cycle in $\mathcal{K}_{J}$. Then the loop
$\gamma_{\lambda}=t(i_{1})\cdot\ell(\lambda)\cdot t(i_{1})^{-1}:\underline{1}\leadsto a_{i_{1}}\leadsto a_{i_{1}}\leadsto\underline{1}$
defines the element in $\pi_{1}(X,\underline{1})$ represented by the word
$R_{\underline{g}}(\lambda,J)=\prod_{t=1}^{k}L_{\underline{g}}(i_{t+1},J\setminus i_{t})\cdot L_{\underline{g}}(i_{t},J\setminus i_{t+1})^{-1}.$
Proof.
The loop $\gamma_{\lambda}$ is the concatenation of loops $t(i_{1})\ell(i_{1},i_{2})\ell(i_{2},i_{3})\dots\ell(i_{k},i_{1})t(i_{1})^{-1}$. By adding the trivial paths $t(i_{s})^{-1}t(i_{s})$ from $a_{i_{s}}$ to $\underline{1}$ and back, we see that $\gamma_{\lambda}$ is homotopic to the loop $\gamma^{\prime}_{\lambda}=t(i_{1})\ell(i_{1},i_{2})t(i_{2})^{-1}\cdot t(i_{2})\ell(i_{2},i_{3})t(i_{3})^{-1}\cdot\dotso\cdot t(i_{k})\ell(i_{k},i_{1})t(i_{1})^{-1}.$ This is the concatenation of loops identified in Lemma 5.17. ∎
Lemma 5.19.
Let $Q\subset\prod_{i=1}^{m}G_{i}$ be an admissible subset, and let ${\underline{g}}\in Q$, $J=\operatorname{supp}{\underline{g}},$ $i\in J$. Then the words $L_{\underline{g}}(i,J)$ and $\operatorname{Red}(L_{\underline{g}}(i,J))$ define the same element in $\pi_{1}(B_{Q},\underline{1}).$
Proof.
We follow the algorithm from Proposition 4.7 which computes $\operatorname{Red}(L_{\underline{g}}(i,J))$. On each step we have a concatenation of loops in $\pi_{1}(B_{Q},\underline{1})$, one of which corresponds to the generator $L_{\underline{g}}(i,J)$, and we replace it with the concatenation of loops $L_{\underline{g}}(j,J)\cdot L_{\underline{g}}(i,J\setminus j)\cdot L_{\underline{g}}(j,J\setminus i)^{-1}$ for some $\{i,j\}\in\mathcal{K}_{J}.$ It is sufficient to show that these two loops are equal in $\pi_{1}(B_{Q},\underline{1})$; equivalently, that the relation (4.1) holds for the corresponding paths in $\pi_{1}(B_{Q},\underline{1})$. The left hand side of this relation is represented by the path $\underline{1}\leadsto v_{\varnothing}\leadsto v_{\{i\}}\leadsto\underline{1}\leadsto v_{\{i\}}\leadsto v_{\{i,j\}}\leadsto\underline{1}.$ This loop is homotopic in $B_{Q}$ to the loop $\underline{1}\leadsto v_{\varnothing}\leadsto v_{\{i\}}\leadsto v_{\{i,j\}}\leadsto\underline{1},$ and the paths $v_{\varnothing}\leadsto v_{\{i\}}\leadsto v_{\{i,j\}}$ and $v_{\varnothing}\leadsto v_{\{j\}}\leadsto v_{\{i,j\}}$ are homotopic since $B_{Q}$ contains the four squares
$\prod_{k\neq i,j}\{g,k\}\times([0_{i},1_{i}]\cup[0_{i},g_{i}])\times([0_{j},1_{j}]\cup[0_{j},g_{j}]).$
Repeating the argument, we obtain the right hand side of this relation. ∎
Corollary 5.20.
The loop $t(\max(J))\cdot s_{i}\cdot t(i)^{-1}\in\pi_{1}(B_{Q},\underline{1})$ represents the word $L_{\underline{g}}(i,J)$ whenever $i\in J$.
The loop $t(i_{1})\ell(i_{1},i_{2})\dots\ell(i_{k},i_{1})t(i_{1})^{-1}\in\pi_{1}(X,\underline{1})$ represents the word $\operatorname{Red}(R_{\underline{g}}(\lambda,J))$ whenever $\lambda=(i_{1},\dots,i_{k},i_{1})$ is a simplicial cycle in $\mathcal{K}_{J}$.∎
We have $B_{Q}=X\cup{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$, where ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$ is contractible and $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}=:C=\iota_{\underline{g}}(|\mathcal{N}(J)|)\simeq|\mathcal{K}_{J}|$ by Proposition 5.9. The path components of $|\mathcal{K}_{J}|$ are represented by the set $\Theta(J)\sqcup\{\max(J)\}$, where the last path component is distinguished; under the homotopy equivalence $X\cap{\mathrm{cc}}(\mathcal{K})_{\underline{g}}\simeq|\mathcal{K}_{J}|$, vertices $i\in J$ correspond to points $a_{i}\in{\mathrm{cc}}(\mathcal{K})_{\underline{g}}$, while edges $\{i,j\}\in\mathcal{K}_{J}$ correspond to paths $a_{i}\leadsto u_{ij}\leadsto a_{j}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$. We denote $m^{\prime}:=\max(J)$, and consider the following paths:
  • For each $i\in\Theta(J)\sqcup\{m^{\prime}\}$, the path $t_{i}:\underline{1}\overset{t(i)}{\leadsto}a_{i}$ in $X$;
  • For each $i\in\Theta(J)$, the path $s_{i}:a_{m^{\prime}}\leadsto v_{\varnothing}\leadsto a_{i}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$;
  • For each cycle $\lambda=(i_{1},\dots,i_{k},i_{k+1}=i_{1})$ in $\mathcal{K}_{J}$, the path $\ell(\lambda):a_{i_{1}}\overset{\ell(i_{1},i_{2})}{\leadsto}a_{i_{2}}\leadsto\dots\leadsto a_{i_{k}}\overset{\ell(i_{k},i_{1})}{\leadsto}a_{i_{1}}$ in ${\mathrm{cc}}(\mathcal{K})_{\underline{g}}$.
This set of paths satisfies the conditions of Lemma 5.1 (here $PC=\Theta(J)\sqcup\{m^{\prime}\}$ and $x=\underline{1}$). Then Corollary 5.20 identify the resulting generators and relations in terms of words in $\operatorname{Cart}(\underline{G},\mathcal{K})$. This completes the proof of Theorem 5.14. ∎
5.7. Infinite case
Recall that a collection of spaces $\{X_{\alpha}\}_{\alpha\in I}$ is directed if for any $\alpha_{1},\alpha_{2}\in I$ there exists $\alpha\in I$ such that $X_{\alpha_{1}}\cup X_{\alpha_{2}}\subset X_{\alpha}.$ In the next lemma, for a presentation $G=\langle\Gamma\mid R\rangle$ we have $\Gamma\subset G$ and $R\subset F(\Gamma)$.
Lemma 5.21.
Let $(X,x_{0})$ be a pointed CW-complex and $\{X_{\alpha}\}$ be a directed set of CW-subcomplexes $X_{\alpha}\subset X$ such that $X=\bigcup_{\alpha}X_{\alpha},$ $x_{0}\in\bigcap_{\alpha}X_{\alpha}$.
Suppose that there are presentations $\pi_{1}(X_{\alpha},x_{0})=\langle\Gamma_{\alpha}\mid R_{\alpha}\rangle$ such that for each inclusion $X_{\alpha}\subset X_{\beta}$ the homomorphism $i_{\alpha\beta}:\pi_{1}(X_{\alpha},x_{0})\to\pi_{1}(X_{\beta},x_{0})$ induces inclusions of presentations. (In other words, $i_{\alpha\beta}$ restricts to an inclusion $\Gamma_{\alpha}\hookrightarrow\Gamma_{\beta}$, and the induced map $(i_{\alpha\beta})_{*}:F(\Gamma_{\alpha})\to F(\Gamma_{\beta})$ restricts to an inclusion $R_{\alpha}\hookrightarrow R_{\beta}$.)
Then $\pi_{1}(X)\cong\Big{\langle}\bigcup_{\alpha}\Gamma_{\alpha}\Big{|}\bigcup_{\alpha}R_{\alpha}\Big{\rangle}.$
Proof.
By [Hat02, Propositions A.4, A.5], every point $x\in X$ has a contractible open neighbourhood $N(x)\subset X$ such that each CW-subcomplex $X_{\alpha}$ is a strict deformation retract of its neighbourhood $V_{\alpha}:=\bigcup_{x\in X_{\alpha}}N(x).$ Replacing $X_{\alpha}$ with $V_{\alpha}$, we now assume that $X_{\alpha}$ are open instead of being CW-subcomplexes.
Moreover, every compact subset $K\subset X$ belongs to some $X_{\alpha(K)}.$ Indeed, we have an open covering $K\subset\bigcup_{\alpha}X_{\alpha};$ choose a finite subcovering $K\subset X_{\alpha_{1}}\cup\dots\cup X_{\alpha_{n}}.$ Since $\{X_{\alpha}\}$ is directed, we have $X_{\alpha_{1}},\dots,X_{\alpha_{n}}\subset X_{\alpha}$ for some $\alpha=\alpha(K).$
Now let $\gamma:(S^{1},{\mathrm{pt}})\to(X,x_{0})$ represent an element $[\gamma]\in\pi_{1}(X,x_{0}).$ Then $\operatorname{Im}\gamma\subset X$ is compact, so $\operatorname{Im}\gamma\subset X_{\alpha}$ for some $\alpha,$ hence $[\gamma]$ belongs to the subgroup of $\pi_{1}(X)$ generated by $\Gamma_{\alpha}.$ It follows that $\pi_{1}(X)$ is generated by the set $\Gamma:=\bigcup_{\alpha}\Gamma_{\alpha}\subset\pi_{1}(X),$ i.e. the natural map $F(\Gamma)\to\pi_{1}(X)$ is surjective.
Similarly, let $R\in\operatorname{Ker}(F(\Gamma)\to\pi_{1}(X))$ be a relation. It depends only on a finite set of generators $[\gamma_{1}],\dots,[\gamma_{N}]\in\Gamma.$ For every $j=1,\dots,N$ we have $[\gamma_{j}]\in\Gamma_{\alpha_{j}}\subset\pi_{1}(X_{\alpha_{j}})$ for some $\alpha_{j}.$ The relation $R$ corresponds to a homotopy between a loop in $X$ and the trivial map. The image of homotopy is an image of unit square, so it is compact and is contained in some $V_{\alpha_{0}}.$ For some $\alpha$, we have $V_{\alpha_{0}},V_{\alpha_{1}},\dots,V_{\alpha_{N}}\subset V_{\alpha}.$
It follows that the word $R$ represents is a relation between $[\gamma_{1}],\dots,[\gamma_{N}]$ in $\pi_{1}(X_{\alpha}),$ hence $R$ belongs to the normal closure of the subset $R_{\alpha}\subset F(\Gamma_{\alpha}).$ It follows that $R$ belongs to the normal closure of $\bigcup_{\alpha}R_{\alpha}\subset F(\bigcup_{\alpha}\Gamma_{\alpha}).$
Proof of Theorem 4.11.
If all groups $G_{i}$ are finite, it suffices to apply Theorem 5.14 to the admissible subset $Q:=\prod_{i=1}^{m}G_{i}.$ In the general case, we apply Lemma 5.21 to the directed covering $\{B_{Q}:Q\subset\prod_{i=1}^{m}G_{i}\text{ is admissible}\}$ of $(B,\underline{1})$ and the presentations of $\pi_{1}(B_{Q},\underline{1})$ provided by Theorem 5.14. (This is a directed covering by Lemma 5.7.) Since any ${\underline{g}}\in\prod_{i=1}^{m}G_{i}$ belongs to an admissible subset, we obtain the set
$\{\widehat{L}_{\underline{g}}(i,\operatorname{supp}{\underline{g}}):{\underline{g}}\in\prod_{i=1}^{m}G_{i},~{}i\in\Theta(\operatorname{supp}{\underline{g}})\}=\{\widehat{L}_{\underline{g}}(i,J):J\subset[m],~{}{\underline{g}}\in\prod_{j\in J}G_{j}^{*},~{}i\in\Theta(J)\}$
of generators for $\operatorname{Cart}(\underline{G},\mathcal{K})$, and similarly with the relations. ∎
6. Discussion
6.1. The case of infinite graphs
The graph product of groups can be defined for any collection of groups $\underline{G}=\{G_{i}:i\in V\}$ and any simple graph $\Gamma\subset\binom{V}{2}$. However, throughout the paper we assumed that $V=[m]$ is a finite set. We claim that Theorem 4.11 holds without this assumption. The required changes in the statement and the proof are listed below.
  • We fix a linear order on $V$. In all formulas, $J\subset V$ is assumed to be finite. Hence the elements $\max(J)\in V$ are well defined and $\mathcal{K}_{J}$ are finite simplicial complexes.
  • In the definition of simplicial complexes (and hence in the definition of the clique complex $\mathcal{K}(\Gamma)$), we require that all faces $I\in\mathcal{K}$ are finite.
  • As in [KM16] and [DK15], infinite polyhedral products of pointed topological pairs $(\underline{X},\underline{A})=\{(X_{i},A_{i},\ast_{i}),i\in V\}$ are defined as unions
    $(\underline{X},\underline{A})^{\mathcal{K}}:=\bigcup_{I\in\mathcal{K}}\bigcup_{\begin{smallmatrix}J\subset V\setminus I,\\ |J|<\infty\end{smallmatrix}}\Big{(}\prod_{i\in I}X_{i}\times\prod_{j\in J}A_{j}\times\prod_{i\notin I\sqcup J}\ast_{i}\Big{)}\subset\prod_{i\in V}X_{i},$
    and are endowed with the colimit topology. Equivalently, consider the restricted direct product
    $\bigoplus_{i\in I}X_{i}:=\{\underline{x}\in\prod_{i\in I}X_{i}:~{}\{i\in I:x_{i}\neq\ast_{i}\}\text{ is finite}\}.$
    Then we define $(\underline{X},\underline{A})^{\mathcal{K}}:=\bigcup_{I\in\mathcal{K}}\Big{(}\prod_{i\in I}X_{i}\times\bigoplus_{j\in V\setminus I}A_{j}\Big{)}\subset\bigoplus_{i\in V}X_{i}.$
  • As noted in [DK15], the Davis’ approach to the proof of Proposition 2.5 (see [Dav12, Theorem 2.18 + Theorem 2.22]) works in the infinite case. (The categorical tools from [PRV04], used by Panov and Veryovkin, also may be applicable here.)
  • In all arguments of Section 5, replace the direct product $\prod_{i=1}^{m}G_{i}$ with the restricted direct product $\bigoplus_{i\in V}G_{i}=\{{\underline{g}}\in\prod_{i\in V}G_{i}:\operatorname{supp}{\underline{g}}\text{ is finite}\}$. The paths $p(I)$ are then defined as infinite concatenations, in which only finite number of paths are non-constant. After omitting the constant subpaths, the concatenation is well defined.
6.2. Number of relations in Cartesian subgroups of graph products
Remark 6.1.
[…]
$\operatorname{rank}\pi_{\operatorname{ab}}\leq\operatorname{def}\operatorname{Cart}(\underline{G},\mathcal{K})+\operatorname{rank}\operatorname{Cart}(\underline{G},\mathcal{K})\leq\operatorname{rank}\pi;$
they disagree if and only if $\operatorname{rank}\pi_{\operatorname{ab}}<\operatorname{rank}\pi.$ Often these ranks are equal. However, the ranks can differ even if all the groups $\Pi_{1}(\mathcal{K}_{J})$ are abelian: for example,
$21=\operatorname{rank}\mathbb{Z}_{6}=\operatorname{rank}(\mathbb{Z}_{2}\oplus\mathbb{Z}_{3})<\operatorname{rank}(\mathbb{Z}_{2}\ast\mathbb{Z}_{3})=$ 2 .
1111222233334444111122223333444411112222333344445555666677778888999910101010111111111212121213131313
Figure 6. The 1-skeleton of a flag complex with $\operatorname{rank}\pi_{\operatorname{ab}}<\operatorname{rank}\pi.$
The smallest flag complex with $\operatorname{rank}\pi>\operatorname{rank}\pi_{\operatorname{ab}}$ known to the author is the clique complex of the graph from Figure 6. This two-dimensional complex has $13$ vertices and its fundamental group is the trefoil knot group $H=\langle a,b\mid a^{2}=b^{3}\rangle$. Since $\operatorname{rank}H=2$, $\operatorname{rank}H_{\operatorname{ab}}=1$ and $H=\Pi_{1}(\mathcal{K})$ is a free summand of $\pi,$ it follows that $\operatorname{rank}\pi>\operatorname{rank}\pi_{\operatorname{ab}}.$
Problem 6.2.
Find the deficiency of $\mathrm{RC}^{\prime}_{\mathcal{K}}$ for $\mathcal{K}$ as in Figure 6.
Although $\operatorname{rank}\pi$ and $\operatorname{rank}\pi_{\operatorname{ab}}$ can differ significantly, they agree when $m$ is small enough: flag simplicial complexes with small $m$ have simpler homotopical properties compared to arbitrary simplicial complexes. For example, one needs $n+1$ vertices to triangulate an $n$-sphere, but $2n$ vertices for a flag triangulation. Also, these ranks agree for surface triangulations:
Proposition 6.3.
Let $\mathcal{K}$ be a full subcomplex in a flag triangulation of a closed oriented surface. Then $\operatorname{rank}\bigoplus_{J\subset[m]}H_{1}(\mathcal{K}_{J};\mathbb{Z})^{\oplus n_{J}}=\sum_{J\subset[m]}n_{J}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J}).$
Proof.
It is sufficient to show that, for each connected full subcomplex $\mathcal{K}_{J}$ of $\mathcal{K}$, the group $H_{1}(\mathcal{K}_{J};\mathbb{Z})$ is free abelian, and $\operatorname{rank}H_{1}(\mathcal{K}_{J};\mathbb{Z})=\operatorname{rank}\pi_{1}(\mathcal{K}_{J})$.
Indeed: for $|\mathcal{K}_{J}|$ a closed oriented surface, this is well known. Otherwise $|\mathcal{K}_{J}|$ is a proper compact subset of a closed surface with finitely generated fundamental group, hence $\pi_{1}(\mathcal{K}_{J})$ is free by [FZ05, Corollary 7]. (More elementary, one can prove that $\pi_{1}(\mathcal{K}_{J})$ is free by collapsing, in the sense of simply homotopy theory, $\mathcal{K}_{J}$ onto its 1-skeleton.) ∎
Remark 6.4.
Due to Proposition 2.5 and [Dav08, Theorem 10.6.1] (see also [Cai17, Theorem 2.3]), $\mathcal{R}_{\mathcal{K}}$ is an aspherical $3$-manifold if and only if $\mathcal{K}$ is a flag triangulation of $S^{2}.$ In this case we know exact number of relations in a “minimal presentation” of the 3-manifold group $\mathrm{RC}^{\prime}_{\mathcal{K}}=\pi_{1}(\mathcal{R}_{\mathcal{K}})$. Note that $H_{1}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})\simeq\bigoplus_{J\subset[m]}{\widetilde{H}}_{0}(\mathcal{K}_{J};\mathbb{Z})$ is free abelian, so the Poincaré duality gives $H_{1}(\mathcal{R}_{\mathcal{K}};\mathbb{Z})^{*}\simeq H_{2}(\mathcal{R}_{\mathcal{K}};\mathbb{Z}).$ Hence our presentation is balanced: the number of generators equals the number of relations. Existence of a balanced presentation is predicted by [Eps61, Theorem 2.5]. Theorem 4.11 provides a minimal balanced presentation of $\mathrm{RC}^{\prime}_{\mathcal{K}}$ in this case. It would be interesting to check if geometric methods of Grujić [Gru22] give a similar presentation.
6.3. Finite coverings and quotients of real moment-angle complexes
The group {±$\{\pm 1\}^{m}$ acts naturally on the real moment-angle complex $\mathcal{R}_{\mathcal{K}}=([-1,1],\{-1,1\})^{\mathcal{K}}$. Quotients $\mathcal{R}_{\mathcal{K}}/H$ by freely acting subgroups $H\subset\{\pm 1\}^{m}$ are known as real toric spaces. In this way one can construct manifolds which admit interesting geometric structures [WY21, Ero22]; the most important examples are the small covers, the real analogues of quasitoric manifolds [DJ91]. Fundamental groups of real toric spaces $\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)$ are thus intermediate subgroups between right-angled Coxeter groups and their commutator subgroups, $\pi_{1}(\mathcal{R}_{\mathcal{K}}/H)\cong\operatorname{Ker}(\mathrm{RC}_{\mathcal{K}}\to\{\pm 1\}^{m}/H)\supset\mathrm{RC}_{\mathcal{K}}^{\prime}$. Classifying spaces of similar subgroups in the right-angled Artin groups were considered in [GILV18].
Explicit presentations for fundamental groups of small covers were given by Wu and Yu [WY21]; in the three-dimensional case, Grujić [Gru22] described a minimal balanced presentation.
Note that $\mathbb{Z}[1/2]$-homology of real toric spaces [CP20] and $\mathbb{Z}$-homology of small covers [CC21] is known. Also, $\mathcal{R}_{\mathcal{K}}/H$ decomposes as a union of $2^{m}/|H|$ copies of ${\mathrm{cc}}(\mathcal{K})$ by construction. Hence one can argue as in Sections 3 and 5 to obtain small presentations of these groups as well as bounds on their rank and deficiency.
Problem 6.5.
Study the fundamental groups of real toric spaces using the methods of this paper.
The same methods apply for fundamental groups of covering spaces for $\mathcal{R}_{\mathcal{K}}$, which are identified with subgroups in $\mathrm{RC}^{\prime}_{\mathcal{K}}$.
6.4. Similarities with graph products of algebras
Our calculation of relations in the group $\mathrm{RC}_{\mathcal{K}}^{\prime}=\pi_{1}(\mathcal{R}_{\mathcal{K}})$ continues the list of parallel results between the case of “real” and “complex” moment-angle complexes $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}\simeq(\mathbb{R},\mathbb{R}\setminus\{0\})^{\mathcal{K}}$ and $\mathcal{Z}_{\mathcal{K}}:=(D^{2},S^{1})^{\mathcal{K}}\simeq(\mathbb{C},\mathbb{C}\setminus\{0\})^{\mathcal{K}}.$ Note that for flag complexes $\mathcal{R}_{\mathcal{K}}$ is aspherical (hence its homotopy properties are governed by the group $\pi_{1}(\mathcal{R}_{\mathcal{K}})$ theory), while $\mathcal{Z}_{\mathcal{K}}$ is simply connected and coformal [Vyl24, Corollary 6.8] (hence its homotopy properties are closely related to the associative algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$).
Many such results are discussed in [PR24, Introduction + Section 7]. Below we expand this list. (Here $\mathcal{K}$ is a flag simplicial complex on $[m],$ and $\mathbf{k}$ is a principal ideal domain.)
On the “homotopy theory side”, we have:
  1. (1)
    $\operatorname{Tor}^{H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})}_{i}(\mathbf{k},\mathbf{k})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J};\mathbf{k})$ by [Vyl22, Theorem 1.2];
  2. (2)
    $\mathrm{cat}(\mathcal{Z}_{\mathcal{K}})=1+\max_{J\subset[m]}\mathrm{cdim}_{\mathbb{Z}}\,\mathcal{K}_{J}$ by [Vyl22, Theorem 1.3], where $\mathrm{cat}(X)$ is the Lusternik–Schnirelmann category of $X$;
  3. (3)
    The ×$\mathbb{Z}\times\mathbb{Z}_{\geq 0}^{m}$-graded $\mathbf{k}$-algebra $H_{*}(\Omega\mathcal{Z}_{\mathcal{K}};\mathbf{k})$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\text{ homogeneous generators modulo }\sum_{J\subset[m]}\mathrm{gen}\,H_{1}(\mathcal{K}_{J};\mathbf{k})\text{ homogeneous relations }$
    [Vyl24, Theorem 5.5] (or, as $\mathbb{Z}$-graded algebra, modulo $\sum_{n\geq 0}\mathrm{gen}\,\bigoplus_{|J|=n}H_{1}(\mathcal{K}_{J};\mathbf{k})$ homogeneous relations [Vyl24, Theorem 5.6]). This presentation is minimal.
  4. (4)
    More generally, if $\underline{X}=(X_{1},\dots,X_{m})$ are simply connected and $\mathbf{k}$ is a field, then the graded $\mathbf{k}$-algebra $H_{*}(\Omega(C\Omega\underline{X},\Omega\underline{X})^{\mathcal{K}};\mathbf{k})$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})\text{ homogeneous generators modulo }\sum_{J\subset[m]}\widetilde{b}_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\dim{\widetilde{H}}_{*}(\Omega X_{i};\mathbf{k})$
    homogeneous relations [Cai24, Corollary 5.14 + Proposition 4.4].
For a discrete group $G$, we have $G\simeq\Omega BG$ as topological monoids, hence $H_{*}(\Omega BG;\mathbb{Z})=H_{0}(\Omega BG;\mathbb{Z})\cong\mathbb{Z}[G]$ as algebras and $\mathrm{H}_{*}(G;\mathbb{Z})=\operatorname{Tor}^{\mathbb{Z}[G]}_{*}(\mathbb{Z},\mathbb{Z})\cong\operatorname{Tor}^{H_{*}(\Omega BG;\mathbb{Z})}_{*}(\mathbb{Z},\mathbb{Z})$ as graded abelian groups. Hence on the “group theory side” we have
  1. (1)
    $\operatorname{Tor}^{H_{*}(\Omega\mathcal{R}_{\mathcal{K}};\mathbb{Z})}_{i}(\mathbb{Z},\mathbb{Z})\cong\bigoplus_{J\subset[m]}{\widetilde{H}}_{i-1}(\mathcal{K}_{J};\mathbb{Z})$ by Proposition 3.4 and Corollary 2.7;
  2. (2)
    $\mathrm{cat}(\mathcal{R}_{\mathcal{K}})=1+\max_{J\subset[m]}\mathrm{cdim}_{\mathbb{Z}}\,\mathcal{K}_{J}$ by [Vyl22, Proposition 5.12];
  3. (3)
    The group $\pi_{1}(\mathcal{R}_{\mathcal{K}})=\mathrm{RC}_{\mathcal{K}}^{\prime}$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\text{ generators modulo }\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})\text{ relations}$
    (Corollary 1.3). The number of generators is minimal.
  4. (4)
    More generally, if $\underline{G}=(G_{1},\dots,G_{m})$ are finite groups, then the group $\operatorname{Cart}(\underline{G},\mathcal{K})=\pi_{1}((E\underline{G},\underline{G})^{\mathcal{K}})$ can be presented by
    $\sum_{J\subset[m]}\widetilde{b}_{0}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})\text{ generators modulo }\sum_{J\subset[m]}\operatorname{rank}\Pi_{1}(\mathcal{K}_{J})\cdot\prod_{i\in J}\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG_{i};\mathbb{Z})$
    relations by Theorem 1.1. (Here we use that $|G|-1=\operatorname{rank}{\widetilde{H}}_{*}(\Omega BG;\mathbb{Z})$.)
In Remark 3.3 we also discuss the similarity between lower bounds for presentations of groups and of connected graded algebras.
Problem 6.6.
What can be said about the loop homology of the polyhedral product $(C\Omega\underline{X},\Omega\underline{X})^{\mathcal{K}}$ if $\mathcal{K}$ is a flag simplicial complex and the spaces $X_{i}$ are neither simply connected nor aspherical?
Remark 6.7.
The parallel results about $\mathcal{Z}_{\mathcal{K}}$ and $\mathcal{R}_{\mathcal{K}}$ can be possibly unified in terms of Morel–Voevodsky’s motivic homotopy theory, since $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}$ and $\mathcal{Z}_{\mathcal{K}}=(D^{2},S^{1})^{\mathcal{K}}$ are real (resp., complex) Betti realisations of the motivic moment-angle complex $\mathcal{Z}_{\mathcal{K}}^{\mathbb{A}^{1}}:=(\mathbb{A}^{1},\mathbb{G}_{m})^{\mathcal{K}},$ a motivic space recently considered by Hornslien [Hor24]. Moreover, the commutator bracket in $\pi_{1}(\mathcal{R}_{\mathcal{K}})$ and the Whitehead bracket in $\pi_{*}(\mathcal{Z}_{\mathcal{K}})$ seem to be the Betti realisations of the (non-bilinear) “motivic Whitehead bracket” $G_{k}\times G_{\ell}\to G_{k+\ell}$ between the non-abelian groups $G_{n}:=[\Sigma(\mathbb{G}_{m})^{\wedge n},\mathcal{Z}_{\mathcal{K}}^{\mathbb{A}^{1}}]$.